### v1

# Assessing High-Risk Systems: An EU AI Act Verification Framework

Abstract.
A central challenge in implementing the AI Act and other AI-relevant regulations in the EU is the lack of a systematic approach to verify their legal mandates. Recent surveys show that this regulatory ambiguity is perceived as a significant burden, leading to inconsistent readiness across Member States. This paper proposes a comprehensive framework designed to help close this gap by organising compliance verification along two fundamental dimensions: the type of method (controls vs. testing) and the target of assessment (data, model, processes, and final product). Additionally, our framework maps core legal requirements to concrete verification activities, serving as a vital bridge between policymakers and practitioners, and aligning legal text with technical standards and best practices. The proposed approach aims to reduce interpretive uncertainty, promote consistency in assessment practices, and support the alignment of regulatory, ethical, and technical perspectives across the AI lifecycle.
1. Introduction
The rapid advancement of Artificial Intelligence (AI) has created significant opportunities alongside new technical, organisational, and regulatory challenges, intensifying the need for trustworthy and well-governed systems. Recent regulatory initiatives, most notably the EU AI Act (eu_ai_act_2024), introduce a comprehensive set of legal obligations for high-risk AI systems. However, a persistent gap remains between these normative requirements and the technical and organisational means available to demonstrate and verify compliance in practice. Translating high-level principles such as fairness, robustness, and transparency into concrete, measurable testing and assurance activities continues to be challenging, requiring structured methodologies, interoperable assessment frameworks, and sustained collaboration across technical, legal, and ethical domains.
In parallel, uncertainty surrounding the Act’s practical implementation is increasingly recognised as a major challenge. This uncertainty manifests at multiple levels: interpretive uncertainty regarding the meaning and scope of specific legal obligations; operational uncertainty concerning their application in everyday development, deployment, and risk management practices; and procedural uncertainty linked to evolving standards, guidance documents, and the gradual designation of notified bodies (deloitte2024ai; dlapiper2025pause; draghi2024competitiveness; arnal2024implementation; lewis2025regulatorylearning; regulatinguncertainty2025). Together, these layers of uncertainty highlight the need for operational tools capable of translating legal expectations into consistently verifiable technical and procedural activities.
The motivation of this paper is to help bridge this gap between regulators, risk managers, developers, technical testers, and certifiers, who often operate with different vocabularies, processes, and levels of maturity. Establishing a shared operational framework can facilitate communication, reduce duplication of effort, and support coordinated assurance practices across the AI lifecycle.
The paper focuses on high-risk AI systems, which are subject to the most extensive obligations under the Act. Prohibited practices are banned outright, while low-risk systems are primarily subject to limited transparency duties that are largely subsumed within the requirements for high-risk systems. Concentrating on this category allows for a detailed examination of how legal obligations can be decomposed into testable, verifiable, and documentable elements, forming the basis for scalable and auditable compliance approaches.
Against this background, the paper is structured around three guiding questions: how high-level legal obligations can be systematically translated into operational components that can be tested, verified, or documented; which dimensions, artefacts, and assessment methods are needed to support a shared operational language among regulators, developers, auditors, and deployers; and how a unified assessment structure can enhance comparability, traceability, and communication across heterogeneous stakeholders and use cases.
The framework proposed in this paper does not aim to be complete or definitive. Rather, it offers a structured methodology intended to evolve alongside regulatory guidance, standardisation efforts, and advances in evaluation techniques. Its purpose is to provide a coherent starting point that can be progressively refined as institutional capacities, testing infrastructures, and sector-specific practices mature.
In addition to presenting the framework at a conceptual level, the paper demonstrates how it operates in practice through a concrete, real-world use case. By applying the framework to an ongoing high-risk AI system in the automotive domain, the paper illustrates how legal obligations can be mapped to technical and procedural controls across the system lifecycle. Overall, the paper presents a structured mapping of the Act’s obligations to currently available technical and organisational mechanisms, organised along four dimensions—data, model, processes, and final products—each associated with specific controls and testing activities.
2. Background
The AI Act introduces a risk-based framework with four categories: prohibited practices, high-risk systems, limited-risk systems, and minimal-risk systems. High-risk systems, defined in Article 6 and Annex III, concern domains where failures could affect health, safety, or fundamental rights, including critical infrastructure, employment, law enforcement, and justice administration. Providers must comply with extensive lifecycle obligations covering risk management, data quality, documentation, transparency, human oversight, and requirements for accuracy, robustness, and cybersecurity, with compliance demonstrated through internal or third-party conformity assessments depending on the system. Translating these obligations into verifiable technical criteria remains challenging. Concepts such as appropriate accuracy, sufficient robustness, or reasonably foreseeable misuse require contextual interpretation grounded in specific use cases and risk profiles. Dataset quality involves dimensions such as representativeness, label accuracy, and temporal stability, each implying different validation tasks. Robustness spans adversarial resilience, distributional stability, and degradation behaviour, necessitating distinct testing methods. Moreover, the Act’s lifecycle logic requires continuous monitoring, not merely point-in-time certification.
Although the technical community has developed numerous fairness, robustness, and transparency tests, they remain fragmented and only loosely connected to regulatory obligations. Recent initiatives provide partial structure: NIST’s AI RMF (NIST_AI_RMF_2025) addresses governance and measurement; ISO/IEC and IEEE standards offer terminology and risk-management guidance (iso_iec_42001_2023); and CEN-CENELEC JTC 21 (CEN-CENELEC_JTC21_2025) is developing harmonized European standards that, under Article 40, will grant presumption of conformity. This work is ongoing, creating a transitional period in which interim frameworks are needed to orient compliance. Notified bodies for third-party assessments (Article 34) are in the process of being designated, while regulatory sandboxes (Article 57) offer controlled environments to test innovative systems and refine assessment approaches.
Persistent challenges remain: misaligned terminology across legal, ethical, and technical communities; fragmented testing methods lacking integration; limited procedural guidance for assessments; insufficient attention to post-deployment phases; and the absence of standardised approaches enabling comparability across systems. The framework proposed here addresses these gaps by offering a structured bridge between the Act’s normative requirements and practical validation processes (mowla2024ai).
3. Methodology
This section outlines the methodological process used to construct the proposed framework, detailing how requirements were identified, categorized, and systematically linked to regulatory, standardization, and scientific sources. The aim is to ensure that every control and testing mechanism rests on a coherent chain of traceability, from high-level legal principles to concrete, technically verifiable methods.
3.1. Requirements
We identified 11 macro-categories of requirements that collectively capture the key dimensions of AI trustworthiness and compliance. The starting point for this categorization was the seven principles of Trustworthy AI defined by the European Commission’s High-Level Expert Group on Artificial Intelligence (HLEG) in its Ethics Guidelines for Trustworthy AI (eu_trustworthy_ai_2019). These principles, formulated prior to the adoption of the EU AI Act, have long served as a foundational reference for aligning AI development with European ethical and societal values. Although the AI Act introduces a more prescriptive and legally binding framework, the essence of these principles remains embedded throughout the regulation: in some cases as stand-alone articles (e.g., transparency, human oversight), and in others distributed across multiple provisions concerning governance, documentation, and system performance.
To complement these ethical foundations with more operational and procedural dimensions, we added four additional categories corresponding to key obligations introduced by the AI Act: quality management, risk management, technical documentation, and record keeping. These categories address the infrastructural and organizational mechanisms that ensure traceability, consistency, and accountability across the AI lifecycle. Together, the original HLEG principles and these four operational domains constitute the 11 macro-requirements underpinning the proposed framework.
Building upon this foundation, each macro-category was further refined into a set of more granular requirements. These were derived through a close reading of the AI Act’s legal provisions, complemented where necessary by interpretative inference from the broader normative intent of the macro-level requirements. This two-tiered decomposition, from principles to operational requirements, enabled a systematic and traceable linkage between high-level ethical concepts and their concrete technical or procedural manifestations.
Once the granular requirements were defined, we proceeded to map each of them to specific guidelines and methods. This mapping was performed hierarchically: first connecting each requirement to relevant guidance sources, and subsequently identifying testing or control methods capable of empirically or procedurally validating compliance. This structured mapping approach ensures conceptual continuity across levels, from principle to requirement, from requirement to guideline, and from guideline to method, thereby providing a coherent foundation for both qualitative governance assessment and quantitative system testing.
3.2. Sources
The identification of control and testing methods followed a structured, multi-stage approach aimed at ensuring both regulatory relevance and technical completeness. The objective was to derive mechanisms that are not only aligned with current legal and ethical requirements, but also grounded in established standards and empirically validated research.
First, we conducted a systematic examination of binding and quasi-binding regulatory sources, beginning with the EU AI Act (eu_ai_act_2024) and the General Data Protection Regulation (GDPR)  (GDPR2016). These texts define the normative baseline for high-risk AI systems in the European context, providing explicit legal obligations related to transparency, robustness, and accountability. Their provisions were used to delimit the scope of each requirement and to establish traceability between legal articles and technical mechanisms.
Second, we reviewed international standards and authoritative guidance documents issued by globally recognised bodies, such as the EU’s General Purpose AI Code of Practice  (gpai-code-practice-3rd), ISACA’s Advanced Audit in AI Official Review Manual (isaca2025aaia), UNESCO’s Recommendation on the Ethics of Artificial Intelligence (unesco2021ethics), ISO/IEC 42001:2023 (iso_iec_42001_2023), ISO23894 (ISO23894) and ISO9001 (ISO9001). These documents were selected because they operationalize high-level principles into implementable governance or audit processes and therefore provide an essential bridge between normative and technical domains.
Third, where specific mappings between legal principles and technical procedures were not sufficiently covered by regulatory or standardization sources, we extended the analysis to the scientific literature. We focused on highly recognised and widely cited works in AI governance, robustness testing, and risk management (e.g., (Brundage2020; Basiri2016; Floridi2018; mundhenk2020)). Inclusion criteria combined citation count, influence of the publication venue, and the relevance of the paper to measurable aspects of trustworthiness such as fairness, reliability, or transparency. This ensured that the resulting mappings draw from the most authoritative and empirically supported methodologies currently available.
Through this layered approach, progressing from regulations to standards to scientific evidence, we aimed to guarantee both the normative validity and the technical soundness of the proposed framework. Each subsection that follows in Section 4 corresponds to one Trustworthy AI requirement and enumerates its associated control and testing mechanisms accordingly.
3.3. Dimensions of Assessment
Figure 1. Mapping between the assessment type and the assessment target dimensions.
Before mapping requirements to methods in Section 4, it is necessary to introduce the conceptual structure underpinning the framework. The analysis of an AI system can be meaningfully organised along two complementary dimensions of assessment. The first dimension concerns the type of assessment, The second dimension concerns the target of assessment. Together, these two axes define the space in which assurance, verification, and risk mitigation take place.
Along the type of assessment axis, the framework distinguishes between controls and testing. Controls are primarily ex-ante activities that establish governance, documentation, and procedural safeguards to ensure responsible development before deployment. Testing, by contrast, consists of ex-post evaluations that empirically examine system behaviour, performance, or statistical properties under specified conditions. Neither approach is hierarchical; rather, they are mutually reinforcing. Controls ensure that development processes are auditable and accountable, while testing provides empirical evidence that outcomes align with intended requirements.
Along the target of assessment axis, the framework identifies four lifecycle layers. The data layer captures aspects related to data governance such as the acquisition, composition, and representativeness of data. Model related concerns such as its architecture, explainability, robustness, and performance are covered by the model dimension. The processes layer encompasses organisational and procedural elements of system development and maintenance, including risk and quality management, documentation, traceability, and lifecycle governance. It reflects how the system is produced and managed rather than what it does. Finally, the final product layer focuses on the behaviour of the deployed system, its interaction with users, and its broader societal or environmental implications, where outputs and impacts can be empirically observed and validated.
The intersection of these two axes yields a comprehensive assessment model that connects organisational assurance with empirical verification across all lifecycle stages. This structure enables evaluators to determine, for each obligation under the AI Act, which combination of assessment type and target is most appropriate, ensuring that governance and verification are treated as interdependent components of a coherent assurance continuum.
4. Mapping
This section presents the mapping between the macro-requirements of Trustworthy AI and the corresponding mechanisms that enable their implementation and verification. Each subsection (R1–R11) details one requirement category, identifying the relevant legal provisions, associated guidelines, and the methods through which compliance can be assessed or demonstrated.
To ensure conceptual clarity, each method is classified along the two analytical dimensions described in 3. Type distinguishes between C (Controls) and T (Tests). Target specifies the primary layer of the AI lifecycle to which the method applies: D (Data), M (Models) , P (Processes) , and FP (Final Product).
4.1. R1: Human Agency and Oversight
Article 14 of the AI Act establishes the principle of human agency and oversight as a cornerstone of trustworthy AI. It requires that AI systems be designed and implemented so that they operate under meaningful human control and that ultimate responsibility remains with human operators. In practice, this means that individuals supervising the system must understand its capabilities and limitations, be able to intervene when necessary, and override its decisions to prevent or mitigate risks. The article also reinforces user autonomy by mandating transparency, informed consent, and the possibility to opt out of AI-driven functionalities. Together, these obligations ensure that human judgment and accountability remain central throughout the design, deployment, and use of AI systems.
Requirement Guidelines Methods Type Target
R1.1: Human in the loop AIA (Art. 14.4.d–e) M1.1: Manual override — Confirm that a human operator can take control of the AI system at any time during operation. C P, FP
AIA (Art. 14.4.a–c) M1.2: Supervisory control — Assess whether human operators are able to supervise the AI system and intervene when necessary. C P, FP
ISO/IEC 42001:2023 (Ann.A.3.2) M1.3: Strategic governance — Ensure that organisational oversight mechanisms for AI deployment are defined and documented. C P
R1.2: User Autonomy AIA (Art. 50.1) M1.4: Informed consent — Check that users are provided with clear information about AI usage and are able to give or withhold consent. C FP
ISACA manual (Sec. 3.2, p.137) M1.5: User preferences — Review whether the system allows adaptation of its behaviour based on user-set preferences. C FP
ISACA manual (Sec. 3.2, p.137) M1.6: Opt-in/out — Confirm that users can enable or disable AI functionalities according to their choices. C FP
R1.3: Oversight Mechanisms ISACA manual (Sec. 2.11.1) M1.7: Audit trails — Ensure that system actions and decisions are recorded in tamper-evident logs for traceability. C P
4.2. R2: Technical Robustness and Safety
Technical robustness and safety under the AI Act (Article 15) concern the system’s ability to operate reliably under normal and adverse conditions, to resist manipulation or degradation, and to recover safely from failures. These requirements ensure that AI systems are resilient to risks arising from data drift, adversarial attacks, or component malfunctions, and that their performance remains consistent throughout the lifecycle. They also mandate mechanisms for fallback and uncertainty management, so that any deviation from intended operation can be detected, controlled, and mitigated in a way that preserves both functionality and safety.
Requirement Guidelines Methods Type Target
R2.1: Resilience & Reliability ISACA manual (Sec. 1.4.4) M2.1: Stress testing — Assess whether the system has been exposed to extreme loads or edge cases to identify potential failure points. T P
ISACA manual (Sec. 2.8.6) M2.2: Drift detection — Confirm that mechanisms are in place to monitor data distribution changes and evaluate behaviour under such shifts. T D, P
Basiri et al. (Basiri2016) M2.3: Chaos engineering — Determine whether resilience has been tested through controlled fault injection in production-like environments. T P
AIA (Art. 15.4); ISACA manual (Sec. 2.9.4) M2.4: Fail-over architecture — Ensure redundant components exist to maintain functionality in case of failure. C P, FP
R2.2: Robustness to Attacks Brundage et al. (Brundage2020); ISACA manual (Sec. 2.13.3) M2.5: Red teaming — Examine whether adversarial testing is conducted to uncover vulnerabilities. T P, FP
R2.3: Fallback & Fail-safe HLEG Guidelines; AIA (Art. 14.4.e) M2.6: Human fallback — Check that the system includes a mechanism to transfer control to a human in the event of failure. C P, FP
R2.4: Accuracy & Uncertainty ISACA manual (Sec. 1.4.3) M2.7: Model calibration — Review whether confidence scores are appropriately aligned with model performance. T M
4.3. R3: Privacy and Data Governance
Article 10 of the AI Act establishes specific obligations concerning data governance and data quality for high-risk AI systems, requiring that datasets be relevant, representative, free of errors, and complete to the extent possible. These provisions complement the data protection principles set out in the GDPR, particularly Articles 5, 25, and 32, which enshrine purpose limitation, privacy by design, and security of processing. Together, they ensure that data is collected, processed, and stored in a lawful, fair, and transparent manner. The combined framework mandates both organisational and technical measures—such as encryption, pseudonymisation, and role-based access control—to prevent unauthorised use and ensure accountability for data-related decisions. Effective data governance thus underpins not only compliance with legal obligations but also the technical robustness and trustworthiness of AI systems across their lifecycle.
Requirement Guidelines Methods Type Target
R3.1: Data Minimisation GDPR (Art. 5.1.b–c) M3.1: Purpose limitation — Verify that only data strictly necessary for a defined processing purpose is collected. C D, P
GDPR (Art. 5.1.b) M3.2: Scope restriction — Assess whether processing is confined to explicitly defined contexts. C D, P
R3.2: Privacy-Enhancing GDPR (Art. 25.1–2; Art. 32.1.a) M3.3: Pseudonymisation — Check whether mechanisms are in place to pseudonymise personal data. C D
R3.3: Access & Security GDPR (Art. 32.1.b–d) M3.4: Role-based access control — Ensure access rights are granted based on organisational roles. C P
GDPR (Art. 32.1.a); ISACA manual (Sec. 2.7.3) M3.5: Encryption — Check that data is encrypted in storage and transmission. T D, P
R3.4: Data Provenance ISACA manual (Sec. 2.7.5) M3.6: Lineage tracking — Confirm that data transformations are logged to ensure traceability. C D, P
ISACA manual (Sec. 3.1.2) M3.7: Metadata versioning — Verify dataset versions and schema changes are tracked. C D, P
R3.5: Consent Management GDPR (Art. 7.1–2) M3.8: Consent tracking — Inspect whether user consents are explicitly obtained, logged, and tied to specific purposes. C P, FP
GDPR (Art. 7.3) M3.9: Granular revocation — Confirm that users can revoke permissions without affecting unrelated consents. C FP, P
4.4. R4: Transparency
Transparency is one of the most fundamental requirements for trustworthy and accountable AI. Under the AI Act, AI systems must be designed and documented in a way that allows users, regulators, and auditors to understand their capabilities, limitations, and decision-making logic. Transparency encompasses both technical and communicative aspects: it includes model explainability, data and process traceability, disclosure of AI usage to end users, and interpretability of outputs in accessible language. Together, these mechanisms ensure that AI operations remain observable and interpretable across all stages of their lifecycle, thereby fostering accountability and enabling meaningful human oversight.
Requirement Guidelines Methods Type Target
R4.1: Explainability ISACA manual (Sec. 3.2, p.139) M4.1: Local interpretability — Verify that the system provides human-understandable explanations of model outputs using recognised local explanation techniques (e.g., feature attribution or surrogate models). T D, M, FP
Mundhenk et al. (mundhenk2020) M4.2: Saliency maps — Check if visualization tools highlight relevant input regions. T D, M, FP
R4.2: Model & Data Traceability GPAI CoP (Sec. Transparency) M4.3: Model cards — Ensure that standardized documentation is available and maintained. C M, P
GPAI CoP (Sec. Transparency) M4.4: Datasheets — Assess whether datasheets describing dataset origin and composition are accessible. C D, P
R4.3: Disclosure of AI Use AIA (Art. 50.1) M4.5: AI identity notices — Verify that users are clearly informed when interacting with an AI system. C FP
GPAI CoP; ISACA manual (Sec. 3.2, p.139) M4.6: Disclaimers — Check whether limitations and assumptions are transparently disclosed. C FP
AIA (Art. 13.3.b.III) M4.7: Risk communication — Assess whether the system communicates known risks and uncertainties. C FP
R4.4: Interpretability for End Users AIA (Art. 13.2) M4.8: Understandable language — Confirm that technical terms are explained using plain language. C FP
4.5. R5: Diversity, Non-discrimination and Fairness
Diversity, non-discrimination, and fairness are central to ensuring that AI systems respect fundamental rights and avoid reinforcing structural biases. Articles 9, 10, and 27 of the AI Act require that systems be designed, trained, and deployed in a way that prevents discriminatory outcomes and ensures inclusive access. This includes assessing data representativeness, applying fairness metrics across protected groups, and incorporating participatory design practices that reflect societal diversity. Fairness thus spans both technical and organisational domains, linking data quality, model evaluation, and stakeholder engagement to the ethical and legal accountability of AI systems.
Requirement Guidelines Methods Type Target
R5.1: Fairness in Data ISACA manual (Sec. 3.2, p.138) M5.1: Bias detection — Evaluate whether datasets are examined for representation gaps. T D, P
ISACA manual (Sec. 3.2, p.138) M5.2: Sampling strategies — Confirm that sampling ensures balanced representation. C D, P
R5.2: Fairness in Algorithms ISACA manual (Sec. 3.2, p.138) M5.3: Protected-class fairness metrics — Verify that fairness metrics are applied across protected groups. T D, M, P
R5.3: Inclusive Design AIA (Rec. 132) M5.4: Accessibility — Check that the system supports assistive technologies. C FP
R5.4: Stakeholder Engagement AIA (Rec. 165) M5.5: Participatory design — Assess whether affected users and communities are involved in design and testing. C FP, P
4.6. R6: Societal and Environmental Well-being
Societal and environmental well-being extend the notion of trustworthy AI beyond individual rights, addressing collective and long-term impacts. Articles 27 and 40 of the AI Act require that systems be aligned with broader ethical and sustainability goals, including environmental protection, energy efficiency, and respect for human dignity. This involves monitoring resource consumption, integrating human rights principles into design, and ensuring that ethical oversight bodies are established to guide responsible innovation. These requirements embed sustainability and societal benefit as integral dimensions of AI governance.
Requirement Guidelines Methods Type Target
R6.1: Environmental Impact UNESCO (Rec. 84); ISACA manual (Sec. 1.17.7) M6.1: Energy-use tracking — Verify that the system monitors computational energy consumption and carbon footprint. T P, FP
R6.2: Ethical Alignment AIA (Art. 27.1) M6.2: Rights-based system — Assess whether system design aligns with fundamental rights. C P
UNESCO (Rec. 58) M6.3: Independent ethics committee — Verify that an independent ethics committee or officer is appointed. C P
4.7. R7: Accountability
Accountability ensures that responsibility for AI-related outcomes is clearly defined and traceable across the system’s lifecycle. Articles 12 and 17 of the AI Act establish the duty to maintain documentation, assign responsibilities, and ensure that auditability and redress mechanisms are in place. Accountability mechanisms link technical and organisational controls to transparency, enabling effective oversight and redress when harm occurs. They require clear responsibility allocation, secure logging, version-controlled documentation, and channels for incident reporting, thereby operationalising the principle that accountability cannot be delegated to the machine.
Requirement Guidelines Methods Type Target
R7.1: Responsibility Assignment ISACA manual (Sec. 3.2, p.129) M7.1: RASCI charts — Confirm that responsibility matrices define roles for each lifecycle activity. C P
R7.2: Auditability ISO/IEC 42001:2023 (Ann.B.6.2.8); ISACA manual (Sec. 2.11.1) M7.2: Logging — Verify that detailed, tamper-evident logs of system operations are maintained and reviewed. C P
ISO/IEC 42001:2023 (Ann.B.8.3) M7.3: External reporting interfaces — Ensure that secure portals are available for third-party auditors. C P
R7.3: Incident Reporting Floridi et al. (Floridi2018) M7.4: Redress systems — Check whether users can report issues or contest AI-driven outcomes. C FP, P
R7.4: Documentation Integrity ISO/IEC 42001:2023 (Ann.B.6.2.5); ISACA manual (Sec. 2.11.1) M7.5: Version-controlled records — Assess whether documentation is versioned and traceable over time. C P
4.8. R8: Quality Management
Article 17 of the AI Act requires providers of high-risk AI systems to establish, implement, document, and maintain a quality management system (QMS). The QMS ensures that the organisation applies consistent procedures for design, testing, data management, and post-market monitoring. It provides a structured approach for achieving and maintaining conformity with the requirements of the AI Act, aligning with recognised standards such as ISO 9001 and ISO/IEC 42001. Quality management thereby supports traceability, continuous improvement, and long-term compliance across the AI system lifecycle.
Requirement Guidelines Methods Type Target
T8.1: Quality policy & objectives ISO 9001:2015 M8.1: Quality Policy Declaration - Verify that top management has issued a formal, documented policy defining quality objectives and regulatory compliance commitments for high-risk AI systems under a Quality Management System (QMS). The QMS would determine, implement, and control processes needed for quality management, and maintain documented information to support operation and provide evidence of conformity. C P
T8.2: Documented QMS procedures ISO 9001:2015 M8.2: Control of documented information - Check that all lifecycle stages (design, development, validation, post-market) have controlled and versioned documented information stored in a document management system. C P
T8.3: Internal quality audits ISO 9001:2015 M8.3: Internal Auditing - Plan and perform periodic audits to verify conformity of processes and compliance with the Regulation; record results and corrective actions. C P
T8.4: Continuous improvement ISO 9001:2015 M8.4: Continuous Improvement - Use audit results, performance data, and feedback to identify non-conformities and improvement opportunities; update processes and objectives accordingly. C P
4.9. R9: Risk Management
Article 9 of the AI Act establishes the obligation to implement a continuous and systematic risk management process throughout the lifecycle of high-risk AI systems. This process covers the identification, analysis, evaluation, and control of risks, as well as post-market monitoring and review. It must ensure that residual risks are acceptable relative to the intended use and that appropriate mitigation measures are applied. The methodology used here aligns with a set of horizontal, cross-sectoral standards, including ISO/IEC 23894 (AI risk management), ISO/IEC 42001 (AI management system), ISO 31000 and ISO 31010 (General risk management principles and techniques), as well as data- and robustness-related standards such as ISO/IEC 5259, ISO 8000, ISO/IEC TR 24027, ISO/IEC 24029, and ISO/IEC 24970. In coming days, when harmonized standards under the AI Act, particularly prEN 18228 (AI Risk Management) and prEN 18286 (AI Quality Management System) will be finalized, they will provide the solid structure for conformity assessment and presumption of conformity once a system is categorized as high-risk.
However, it is essential to underline that none of these standards, both international ISO/IEC standards and future harmonized European standards, determine whether an AI system is classified as high-risk under Article 6. Classification is governed exclusively by the legal criteria of the AI Act, particularly Annex I and Annex III. Standards are used in this report solely to support structured analysis, documentation and evidence generation, and they become fully applicable only after an AI system has been classified as high-risk.
In addition, the methodology deliberately excludes sector-specific or application-specific standards, such as functional-safety standards (e.g., IEC 61508 or ISO 13849), medical-device standards (e.g., ISO 14971) or automotive safety standards (e.g., ISO 26262). These frameworks are highly relevant within their respective regulatory regimes but are not part of the horizontal, cross-sector risk-classification process defined by the AI Act and therefore are not included in the present analysis.
Requirement Guidelines Methods Type Target
T9.1: System Definition ISO/IEC 42001 (§§ 4.3, 6.1.2); ISO/IEC 23894 § 5.3 M9.1 Description of AI system – define intended purpose for Art 6 mapping and record system details C P
T9.2: Risk identification ISO/IEC 42001 (§§ 6.1.2, 7.5); ISO/IEC 23894 § 5.4; ISO 31000; ISO 31010; ISO/IEC 5259; ISO 8000; ISO/IEC TR 24027 M9.2 Structured risk identification feeding both Art. 6 classification and Art. 9 risk management C P
T9.3: Risk analysis ISO/IEC 23894; ISO/IEC 42001; ISO 31000; ISO 31010; M9.3 Core classification step – apply AI Act Art. 6: Annex I & Annex III gates based on structured analysis C P
T9.4: Risk evaluation & prioritisation ISO/IEC 23894; ISO/IEC 42001; ISO 31000; ISO 31010 M9.4 Finalise classification decision and justification recording C P
T9.5: Risk control measures ISO/IEC 23894 §§ 5.5–5.7; ISO/IEC 42001; ISO 31000; ISO/IEC 24029-1/-2; ISO/IEC TR 24027; ISO/IEC 5259; ISO 8000; ISO/IEC 27001; 27005 M9.5 Risk control strategy documented and aligned with AI Act essential requirements C FP
T9.6: PMM: Planning ISO/IEC 42001 (§§ 8.4, 9.1, 7.5); ISO/IEC 23894 (§§ 5.3, 5.7); ISO 31000; ISO/IEC 24970; ISO/IEC 27001; 27005 M9.6 PMM system established with structured monitoring requirements C P
T9.7: PMM: Data collection & analysis ISO/IEC 42001 (§§ 9.1, 9.3, 7.5); ISO/IEC 23894 (§§ 5.5–5.7); ISO 31000; ISO/IEC 24970; ISO/IEC 27001; 27005 M9.7 Continuous assessment and early detection of anomalies or new hazards T FP
T9.8: PMM: Review & CAPA ISO/IEC 42001 (§§ 8.3, 10, 7.5); ISO/IEC 23894 (§§ 5.6, 5.8); ISO 31000 M9.8 CAPA implemented, system updated, continuous compliance maintained C P
T9.9: Incident reporting ISO/IEC 42001 (§§ 8.4, 9.1, 7.5); ISO/IEC 23894 (§§ 5.7, 5.8); ISO/IEC 24970; ISO/IEC 27035 M9.9 Compliant incident reporting + full documentation + integration of learnings C P
4.10. R10: Technical Documentation
Article 11 and Annex IV of the AI Act require providers of high-risk AI systems to prepare and maintain comprehensive technical documentation demonstrating compliance with the regulation. This documentation must provide sufficient detail to assess the system’s conformity, including its intended purpose, design specifications, data sources, testing methods, and risk management processes. Proper documentation ensures transparency, enables traceability of design choices, and supports both internal governance and external conformity assessment.
Requirement Guidelines Methods Type Target
R10.1: System overview & intended purpose prEN 18286; ISO 9001; ISO/IEC 42001 M10.1 System overview and intended purpose documented in alignment with Annex IV and QMS requirements. C FP
R10.2: Design specifications & architecture prEN 18286; ISO 9001; ISO/IEC 42001 M10.2 Architecture and design specifications captured with complete traceability to risks and requirements. C M, P
R10.3: Dataset description & provenance prEN 18286; ISO 9001; ISO/IEC 42001 M10.3 Dataset provenance, quality attributes and limitations documented in compliance with Art. 10 and Annex IV. C D
R10.4: Performance metrics & validation prEN 18286; ISO 9001; ISO/IEC 42001 M10.4 Performance, robustness and validation evidence compiled and traceable to requirements, risks and intended purpose. T FP
R10.5: Compliance evidence mapping prEN 18286; ISO 9001; ISO/IEC 42001 M10.5 Structured compliance mapping demonstrating how Annex IV and Article 17 obligations are satisfied. C P
4.11. R11: Record-keeping
Article 18 of the AI Act mandates that providers retain automatically generated logs and relevant records to ensure traceability and accountability throughout the AI lifecycle. Record-keeping supports post-market monitoring, facilitates incident investigation, and provides evidence for conformity assessments. It must include documentation of model versions, dataset lineage, and data retention or deletion policies, consistent with organisational governance and data protection requirements.
Requirement Guidelines Methods Type Target
R11.1: Operational logs prEN 18286; ISO 9001; ISO/IEC 42001 M11.1 Operational logging system in place with records supporting Art. 12 transparency and PMM requirements. C FP
R11.2: Version control records prEN 18286; ISO 9001; ISO/IEC 42001 M11.2 Version control ensures full traceability and reproducibility of AI system states over time. C M
R11.3: Model & dataset lineage prEN 18286; ISO 9001; ISO/IEC 42001 M11.3 Full lineage enables auditing, incident analysis, risk traceability and evidence-based updates. C D, M
R11.4: Retention & deletion policies prEN 18286; ISO 9001; ISO/IEC 42001 M11.4 Retention/deletion policies implemented to ensure compliance, minimise risk and maintain only necessary records. C P
5. Example application
This section illustrates how the proposed framework can be applied in practice through a concrete example drawn from a real, ongoing high-risk AI use case currently being addressed in collaboration with a Swedish automotive manufacturer. While the framework itself is generic and technology-agnostic, this example demonstrates how it can be used to structure assurance activities across the full lifecycle of an AI system, from design and testing to validation and documentation.
The use case concerns an AI-based system for detecting cyberattacks within connected vehicles. Modern vehicles rely on internal communication networks that enable electronic components to exchange information in real time. Protecting these networks against malicious interference is critical, as cyberattacks may compromise both vehicle safety and reliability. To mitigate these risks, automotive manufacturers increasingly deploy Intrusion Detection Systems that monitor in-vehicle communications and identify abnormal or potentially harmful behaviour.
In this setting, the IDS employs AI techniques to analyse network traffic within the vehicle and detect patterns associated with known or emerging attacks. The system processes data from in-vehicle networks, extracts relevant features, and uses machine-learning models to distinguish normal operation from suspicious activity. To support transparency and engineering oversight, the system also integrates explainability mechanisms that allow engineers to understand why specific activities are flagged as intrusions.
Designed for real-time deployment in vehicle control units or central gateways, this real-world use case illustrates how the proposed framework can be applied in a complex, safety-critical automotive context. In particular, the framework enables a structured and traceable assessment of data handling, model behaviour, explainability, and documentation, providing evidence to support safety, cybersecurity, and regulatory requirements.
Legal Requirement (AI Act) Test / Control Type Method Instantiation in the Use Case
R1: Human Agency & Oversight (Art. 14) T1.1 Human-in-the-loop M1.1 Manual override Engineers and operators can override IDS decisions and reclassify labels during development.
M1.2 Supervisory control IDS outputs are logged and verified by safety and security engineers before integration into final decisions.
M1.3 Strategic governance Ensure a documented organisational oversight structure: define who in the team (e.g., vehicle OEM cybersecurity lead) is responsible for authorising deployment and change management of the IDS.
T1.2 User Autonomy M1.4 Informed consent Data collection from vehicles was conducted under user consent for research and model development.
M1.5 User preferences Provide vehicle end-users or clients (e.g., fleet operators) the ability to configure IDS alert sensitivity or disable non-critical automated responses (while still keeping core safety monitoring active).
T1.3 Oversight Mechanisms M1.7 Audit trails The IDS decisions, rule activations, and system overrides are logged for auditability.
R2: Technical Robustness & Safety (Art. 15) T2.1 Resilience & Reliability M2.1 Stress testing Baseline models were evaluated under high-throughput simulated traffic conditions.
M2.2 Drift detection Monitor for changes over time in the in-vehicle network traffic (e.g., ECU updates, different firmware) and detect when feature distributions deviate; trigger retraining or model review.
M2.4 Fail-over architecture Ensure the IDS has a fallback mode: if the neuro-symbolic model fails or confidence drops, revert to a simpler rule-based guard or human-supervised monitoring.
T2.2 Robustness to Attacks M2.5 Red teaming The system was tested with synthetic adversarial attacks, including crafted PTP manipulations.
T2.3 Fallback & Fail-safe M2.6 Human fallback Low-confidence detections trigger alerts requiring human verification.
T2.4 Accuracy & Uncertainty M2.7 Model calibration Performance metrics include confidence intervals; rule activation thresholds are tunable.
R3: Privacy & Data Governance (Art. 10; GDPR) T3.1 Data Minimisation M3.1 Purpose limitation Only relevant features for detection (e.g., timestamp, ID, payload length) are processed.
T3.2 Privacy-Enhancing M3.3 Pseudonymisation When capturing vehicle network traffic for training, any identifiable driver or vehicle identity data is pseudonymised or removed, retaining only features necessary for intrusion detection.
T3.3 Access & Security M3.4 Role-based access control Raw data is accessible only to authorised research engineers; results are pseudonymised.
R4: Transparency (Art. 13, 50) T4.1 Explainability M4.1 SHAP / M4.2 LIME SHAP is used to validate feature contributions; LNN rules provide symbolic reasoning.
T4.2 Model & Data Traceability M4.4 Model cards All models used are accompanied by documentation on training data, assumptions, and risks.
M4.5 Datasheets Provide a datasheet for the dataset used in the IDS: describe origin (e.g., vehicles from the partner), format, attack types covered, limitations, and preprocessing steps to support traceability in audits.
T4.3 Disclosure of AI Use M4.6 AI identity notice In future deployments, HMI interfaces will indicate when AI-based IDS is active.
R5: Fairness & Non-Discrimination (Art. 9, 10) T5.1 Fairness in Data M5.1 Data bias detection Attack data were balanced to ensure representative detection performance.
T5.4 Stakeholder Engagement M5.5 Participatory design review Security experts and domain engineers reviewed and refined rule design and evaluation criteria.
R7: Accountability (Art. 12, 17) T7.1 Responsibility Assignment M7.1 Responsibility-matrix verification Roles for data handling, model design, and validation are clearly documented.
T7.2 Auditability M7.2 Logging All detection events and rule-based decisions are logged and version controlled.
T7.3 Incident Reporting M7.4 Redress systems False positives or missed detections can be flagged through internal validation interfaces.
T7.4 Documentation Integrity M7.5 Version-controlled records Maintain version control for IDS models, rule sets, training datasets, and change logs (who changed what, when, and why) to support traceability and post-incident investigation.
R8: Quality Management (Art. 17) T8.1 Quality policy & objectives M8.1 Quality Policy Declaration The company’s cybersecurity and safety management define a documented QMS policy for AI-based IDS development, specifying objectives on detection accuracy, explainability, robustness, and regulatory alignment; relevant engineers receive QMS training.
T8.2 Documented QMS procedures M8.2 Control of documented information IDS development artifacts (model architectures, LNN rule sets, preprocessing pipelines, datasets, validation reports) are placed under controlled documentation with restricted access and traceability.
T8.3 Internal quality audits M8.3 Internal auditing Periodic internal audits review model lineage logs, robustness testing outcomes, fairness checks, and documentation completeness; non-conformities trigger corrective actions.
T8.4 Continuous improvement M8.4 Continuous improvement Findings from audits and monitoring are used to update models, rules, thresholds, and documentation throughout the lifecycle.
R9: Risk classification, management, and post marketing (Art. 6, Art. 9) T9.1 System Definition M9.1 Description of AI system The automotive AI‑based IDS is defined as a system intended to detect cyber intrusions targeting in‑vehicle networks and connected services.
T9.2 Risk identification M9.2 Structured risk identification Structured risk identification places the automotive IDS as preliminarily high‑risk due to safety‑critical vehicle functions.
T9.3 Risk analysis M9.3 Core classification step The IDS is assessed against AIA Annex I & III, with critical‑infrastructure and safety‑component gating confirming high‑risk applicability.
T9.5 Risk control measures M9.5 Risk control strategy Risk controls are established, covering safety, robustness, explainability, human oversight, and documentation through git platform in automotive contexts.
T9.6 PMM planning M9.6 PMM system established A post‑market monitoring plan is defined to track performance, drift, and safety‑relevant anomalies across vehicle fleets.
R10: Technical Documentation (Art. 11; Annex IV) T10.1 System overview & intended purpose M10.1 Annex-IV compliant description Provide a structured description of the IDS: architecture, model pipeline, LNN rule types, deployment targets (ECUs, gateways), and intended function (early anomaly detection).
T10.2 Design specifications & architecture M10.2 Architecture traceability Document the full dataflow: CAN/Ethernet packet capture $\rightarrow$ preprocessing $\rightarrow$ feature extraction $\rightarrow$ ML/LNN inference $\rightarrow$ explainability layer $\rightarrow$ alert routing, with traceability from risks to design decisions.
T10.3 Dataset description & provenance M10.3 Dataset provenance Record dataset sources (vehicle captures), attack types, collection conditions, preprocessing steps, firmware versions when relevant, and known limitations (e.g., limited coverage of rare timing attacks).
T10.4 Performance metrics & validation M10.4 Validation evidence Provide detection rate, latency under high traffic, robustness under noise, calibration curves, and interpretability validation using SHAP and LNN reasoning.
T10.5 Compliance evidence mapping M10.5 Mapping to obligations Map each AI Act requirement to evidence: oversight logs, stress-test results, explainability artefacts, model cards, dataset datasheets, QMS records, and PMM outputs.
R11: Record-keeping (Art. 12, 18) T11.1 Operational logs M11.1 Operational logging Maintain logs of detections, LNN rule activations, SHAP explanations shown for review, fallback events, and red-team outcomes; store logs securely in automotive backend systems.
T11.2 Version control records M11.2 Version control Track versions of ML models, LNN logic rules, preprocessing code, and dataset snapshots in a versioned repository with approvals and traceable change history.
T11.3 Model & dataset lineage M11.3 Lineage documentation Maintain lineage links from raw CAN/Ethernet captures to preprocessing scripts, training datasets, trained model versions, and deployed artefacts to support reproducibility and audit.
T11.4 Retention & deletion policies M11.4 Retention/deletion Define retention periods for logs, dataset snapshots, and models; ensure GDPR compliance when identifiable metadata appears; enforce secure deletion mechanisms.
6. Discussion and Conclusion
The framework presented in Section 3 operationalises compliance by establishing explicit correspondences between legal requirements, recognised standards, and assessment protocols. Its stratified architecture—distinguishing controls and testing procedures across data, model, process, and product dimensions—articulates how different modes of assurance interrelate throughout the AI system lifecycle. This structure facilitates systematic self-assessment by delineating which technical and organisational elements require documented evidence of conformity, thereby supporting the structuring of documentation and monitoring activities. Such guidance may prove particularly valuable for organisations lacking established compliance infrastructure.
Within regulatory sandboxes, the framework serves as a common reference architecture for aligning testing protocols and reporting practices with AI Act requirements. By translating legal principles into empirically verifiable criteria, it enables more transparent and structured dialogue between developers and regulatory authorities. Furthermore, it promotes comparability across sandbox experiments and supports coherent assessment of compliance readiness at different stages of system development. For formal conformity assessment, the mapping provides auditors and notified bodies with an instrument for distinguishing procedural from empirical evidence, enhancing traceability, and complementing sector-specific standards through a unified vocabulary for evaluating regulatory conformity.
Beyond immediate applications, the framework contributes to the longer-term objective of partially or fully automated compliance verification. By formalising relationships between legal requirements, control mechanisms, and testing procedures, it establishes a foundation for computational systems capable of selecting, executing, and interpreting appropriate assessment protocols. This aligns with emerging initiatives that translate regulatory obligations into executable testing workflows and interactive monitoring dashboards (buscemi2025sandbox; ki_prueftoolsuite_2025). A natural extension involves developing an ontology that links each legal provision to its associated controls and testable mechanisms, thereby supporting partial automation and interoperability across assessment tools and regulatory databases. Incorporating a temporal dimension would additionally enable longitudinal analysis of how regulatory frameworks, assessment methodologies, and compliance expectations evolve, facilitating both retrospective evaluation and prospective governance planning.
The framework possesses inherent limitations. Any mapping from legal provisions to technical methods necessarily involves interpretive judgement, and the appropriateness of individual assessment procedures remains context-dependent. The deliberately generic level of abstraction may constrain direct application without domain-specific adaptation. Consequently, the framework should be understood not as a prescriptive checklist but as a reference architecture subject to iterative refinement as standards mature and regulatory practice develops. Future research directions include developing prioritisation schemes among requirements, integrating continuous monitoring mechanisms, and extending the framework to encompass additional EU instruments relevant to AI governance.
This study provides an initial architectural foundation for decomposing AI Act obligations into structured, empirically testable components. By establishing coherent mappings among requirements, assessment methods, and implementation guidelines, it seeks to bridge epistemic and operational divides among regulators, risk managers, developers, technical auditors, and prospective certification bodies. The framework is conceived as an evolving infrastructure—supporting transparent, consistent, and traceable compliance verification while fostering evidence-based dialogue between technical and legal communities. It represents both a methodological contribution and an invitation to collaborative development toward a shared, adaptive, and continuously improving ecosystem for trustworthy AI assurance.

### v2

@@
-Assessing High-Risk Systems: An EU AI Act Verification Framework
+Assessing High-Risk AI Systems under the EU AI Act: From Legal Requirements to Technical Verification
 Abstract.
-A central challenge in implementing the AI Act and other AI-relevant regulations in the EU is the lack of a systematic approach to verify their legal mandates.
-Recent surveys show that this regulatory ambiguity is perceived as a significant burden, leading to inconsistent readiness across Member States.
-This paper proposes a comprehensive framework designed to help close this gap by organising compliance verification along two fundamental dimensions: the type of method (controls vs.
-testing) and the target of assessment (data, model, processes, and final product).
-Additionally, our framework maps core legal requirements to concrete verification activities, serving as a vital bridge between policymakers and practitioners, and aligning legal text with technical standards and best practices.
-The proposed approach aims to reduce interpretive uncertainty, promote consistency in assessment practices, and support the alignment of regulatory, ethical, and technical perspectives across the AI lifecycle.
+The implementation of the AI Act requires practical mechanisms to verify compliance with legal obligations, yet concrete and operational mappings from high-level requirements to verifiable assessment activities remain limited, contributing to uneven readiness across Member States.
+This paper presents a structured mapping that translates high-level AI Act requirements into concrete, implementable verification activities applicable across the AI lifecycle.
+The mapping is derived through a systematic process in which legal requirements are decomposed into operational sub-requirements and grounded in authoritative standards and recognised practices.
+From this basis, verification activities are identified and characterised along two dimensions: the type of verification performed and the lifecycle target to which it applies.
+By making explicit the link between regulatory intent and technical and organisational assurance practices, the proposed mapping reduces interpretive uncertainty and provides a reusable reference for consistent, technology-agnostic compliance verification under the AI Act.
 1. Introduction
-The rapid advancement of Artificial Intelligence (AI) has created significant opportunities alongside new technical, organisational, and regulatory challenges, intensifying the need for trustworthy and well-governed systems.
-Recent regulatory initiatives, most notably the EU AI Act (eu_ai_act_2024), introduce a comprehensive set of legal obligations for high-risk AI systems.
-However, a persistent gap remains between these normative requirements and the technical and organisational means available to demonstrate and verify compliance in practice.
-Translating high-level principles such as fairness, robustness, and transparency into concrete, measurable testing and assurance activities continues to be challenging, requiring structured methodologies, interoperable assessment frameworks, and sustained collaboration across technical, legal, and ethical domains.
-In parallel, uncertainty surrounding the Act’s practical implementation is increasingly recognised as a major challenge.
-This uncertainty manifests at multiple levels: interpretive uncertainty regarding the meaning and scope of specific legal obligations; operational uncertainty concerning their application in everyday development, deployment, and risk management practices; and procedural uncertainty linked to evolving standards, guidance documents, and the gradual designation of notified bodies (deloitte2024ai; dlapiper2025pause; draghi2024competitiveness; arnal2024implementation; lewis2025regulatorylearning; regulatinguncertainty2025).
-Together, these layers of uncertainty highlight the need for operational tools capable of translating legal expectations into consistently verifiable technical and procedural activities.
-The motivation of this paper is to help bridge this gap between regulators, risk managers, developers, technical testers, and certifiers, who often operate with different vocabularies, processes, and levels of maturity.
-Establishing a shared operational framework can facilitate communication, reduce duplication of effort, and support coordinated assurance practices across the AI lifecycle.
-The paper focuses on high-risk AI systems, which are subject to the most extensive obligations under the Act.
-Prohibited practices are banned outright, while low-risk systems are primarily subject to limited transparency duties that are largely subsumed within the requirements for high-risk systems.
-Concentrating on this category allows for a detailed examination of how legal obligations can be decomposed into testable, verifiable, and documentable elements, forming the basis for scalable and auditable compliance approaches.
-Against this background, the paper is structured around three guiding questions: how high-level legal obligations can be systematically translated into operational components that can be tested, verified, or documented; which dimensions, artefacts, and assessment methods are needed to support a shared operational language among regulators, developers, auditors, and deployers; and how a unified assessment structure can enhance comparability, traceability, and communication across heterogeneous stakeholders and use cases.
-The framework proposed in this paper does not aim to be complete or definitive.
-Rather, it offers a structured methodology intended to evolve alongside regulatory guidance, standardisation efforts, and advances in evaluation techniques.
-Its purpose is to provide a coherent starting point that can be progressively refined as institutional capacities, testing infrastructures, and sector-specific practices mature.
-In addition to presenting the framework at a conceptual level, the paper demonstrates how it operates in practice through a concrete, real-world use case.
-By applying the framework to an ongoing high-risk AI system in the automotive domain, the paper illustrates how legal obligations can be mapped to technical and procedural controls across the system lifecycle.
-Overall, the paper presents a structured mapping of the Act’s obligations to currently available technical and organisational mechanisms, organised along four dimensions—data, model, processes, and final products—each associated with specific controls and testing activities.
-2. Background
-The AI Act introduces a risk-based framework with four categories: prohibited practices, high-risk systems, limited-risk systems, and minimal-risk systems.
-High-risk systems, defined in Article 6 and Annex III, concern domains where failures could affect health, safety, or fundamental rights, including critical infrastructure, employment, law enforcement, and justice administration.
-Providers must comply with extensive lifecycle obligations covering risk management, data quality, documentation, transparency, human oversight, and requirements for accuracy, robustness, and cybersecurity, with compliance demonstrated through internal or third-party conformity assessments depending on the system.
-Translating these obligations into verifiable technical criteria remains challenging.
-Concepts such as appropriate accuracy, sufficient robustness, or reasonably foreseeable misuse require contextual interpretation grounded in specific use cases and risk profiles.
-Dataset quality involves dimensions such as representativeness, label accuracy, and temporal stability, each implying different validation tasks.
-Robustness spans adversarial resilience, distributional stability, and degradation behaviour, necessitating distinct testing methods.
-Moreover, the Act’s lifecycle logic requires continuous monitoring, not merely point-in-time certification.
-Although the technical community has developed numerous fairness, robustness, and transparency tests, they remain fragmented and only loosely connected to regulatory obligations.
-Recent initiatives provide partial structure: NIST’s AI RMF (NIST_AI_RMF_2025) addresses governance and measurement; ISO/IEC and IEEE standards offer terminology and risk-management guidance (iso_iec_42001_2023); and CEN-CENELEC JTC 21 (CEN-CENELEC_JTC21_2025) is developing harmonized European standards that, under Article 40, will grant presumption of conformity.
-This work is ongoing, creating a transitional period in which interim frameworks are needed to orient compliance.
-Notified bodies for third-party assessments (Article 34) are in the process of being designated, while regulatory sandboxes (Article 57) offer controlled environments to test innovative systems and refine assessment approaches.
-Persistent challenges remain: misaligned terminology across legal, ethical, and technical communities; fragmented testing methods lacking integration; limited procedural guidance for assessments; insufficient attention to post-deployment phases; and the absence of standardised approaches enabling comparability across systems.
-The framework proposed here addresses these gaps by offering a structured bridge between the Act’s normative requirements and practical validation processes (mowla2024ai).
+Progress in AI has created substantial opportunities while also introducing significant technical, organisational, and regulatory challenges, particularly as AI systems are increasingly deployed in critical domains where failures may have serious societal consequences (bengio2024managing; wang2022artificial).
+In response, regulatory efforts worldwide are intensifying to promote trustworthiness in AI development and deployment (huang2024; smuha2021race; novelli2024taking).
+Within this context, the EU Artificial Intelligence Act (AI Act) (eu_ai_act_2024) represents one of the most comprehensive regulatory responses to date, establishing a risk-based framework that mandates ex ante verification and ongoing oversight for high-risk AI systems.
+Despite this progress, a persistent gap remains between high-level legal requirements and the concrete verification activities needed to demonstrate compliance in practice.
+This gap reflects the absence of explicit and shared mappings between regulatory obligations and implementable assurance practices, resulting in fragmented compliance efforts across stakeholders and Member States and limiting comparability and regulatory learning.
+In practice, organisations often struggle to translate abstract legal obligations into verifiable controls, tests, and documentation artefacts that can be applied consistently across the AI lifecycle.
+These challenges are compounded by three interrelated forms of uncertainty.
+Interpretive uncertainty concerns the contextual meaning and scope of legal obligations; operational uncertainty relates to how such obligations can be instantiated across heterogeneous development and deployment practices; and procedural uncertainty arises from evolving standards, guidance documents, and the gradual designation of notified bodies (deloitte2024ai; dlapiper2025pause; draghi2024competitiveness; arnal2024implementation; lewis2025regulatorylearning; regulatinguncertainty2025).
+These uncertainties are further amplified by the emergent behaviour of modern generative AI systems, whose properties arise from complex interactions rather than explicit programming (holtzmanGenerativeModelsComplex2025; koltLessonsComplexSystems2025a).
+As a horizontal regulation relying on harmonised standards, post-market monitoring, and iterative guidance, the AI Act embodies a model of regulatory learning in which compliance expectations are progressively refined through implementation experience and verification outcomes (zakiConceptualisingOrganisationalPolicy2025; lewis2025regulatorylearning; schrepelAdaptiveRegulation2025).
+Such learning depends on the availability of structured, comparable, and traceable verification evidence, which remains fragmented across actors, disciplines, and use cases.
+This paper addresses the need for an explicit and operational mapping between AI Act requirements and concrete verification activities that can be executed, documented, and compared across the AI lifecycle.
+The objective is to operationalise compliance by making the link between regulatory intent and technical and organisational assurance practices explicit, reproducible, and auditable.
+The focus is on high-risk AI systems as defined in Article 6 and Annex III, which concern application domains where failures may adversely affect health, safety, or fundamental rights.
+Compliance must be demonstrated through conformity verification procedures under Articles 43 and 44, either via internal control procedures (Annex VI) or third-party assessment by notified bodies (Annex VII), both requiring verifiable evidence spanning design, development, deployment, and post-market monitoring.
+Translating these obligations into concrete assessment activities remains challenging.
+Many provisions of the AI Act are inherently context-dependent, requiring case-specific interpretation grounded in the system’s intended use, operational environment, and evolving risk profile.
+Requirements related to data governance, robustness, accuracy etc.
+encompass multiple dimensions, each necessitating distinct validation and testing methodologies, while the lifecycle-oriented nature of the Act entails continuous monitoring and post-deployment evaluation rather than reliance on point-in-time certification alone.
+Existing initiatives provide partial foundations for AI verification, but none offers a complete and operational mapping between legal requirements and verification activities.
+The NIST AI Risk Management Framework (NIST_AI_RMF_2025) provides high-level guidance but is not aligned with EU-specific legal obligations, while ISO/IEC standards such as ISO/IEC 42001 (iso_iec_42001_2023) establish process-oriented management requirements without specifying AI Act–specific verification procedures.
+European standardisation efforts within CEN–CENELEC JTC 21 (CEN-CENELEC_JTC21_2025) aim to develop harmonised standards granting presumption of conformity under Article 40, but this work remains ongoing.
+Against this background, the central contribution of this paper is a structured and operational mapping that translates high-level AI Act requirements into concrete verification activities applicable across the AI lifecycle.
+The mapping is constructed through a systematic decomposition of legal obligations into operational sub-requirements and is grounded in authoritative standards and recognised engineering and governance practices.
+To ensure consistency and comparability, verification activities are characterised along two dimensions capturing both verification type and lifecycle target.
+The applicability of the mapping is illustrated through a real-world case study involving a high-risk AI system in the automotive domain.
+Within this scope, the paper addresses three guiding questions that structure the construction of the proposed mapping: (i) how high-level legal obligations can be decomposed into concrete verification activities; (ii) which verification dimensions are required to ensure comparability and traceability across heterogeneous actors; and (iii) how a shared verification structure can support communication and regulatory learning across use cases and sectors.
+While related efforts exist (e.g.
+(hernandez2025open)), the proposed mapping provides a higher level of granularity and operationalisability.
+It is not intended to be prescriptive or exhaustive, but to serve as a reusable reference that can evolve alongside regulatory guidance, standardisation efforts, and advances in evaluation techniques, thereby supporting more consistent conformity verification and evidence-based regulatory learning over time.
 2. Methodology
-This section outlines the methodological process used to construct the proposed framework, detailing how requirements were identified, categorized, and systematically linked to regulatory, standardization, and scientific sources.
-The aim is to ensure that every control and testing mechanism rests on a coherent chain of traceability, from high-level legal principles to concrete, technically verifiable methods.
-3.1. Requirements
-We identified 11 macro-categories of requirements that collectively capture the key dimensions of AI trustworthiness and compliance.
-The starting point for this categorization was the seven principles of Trustworthy AI defined by the European Commission’s High-Level Expert Group on Artificial Intelligence (HLEG) in its Ethics Guidelines for Trustworthy AI (eu_trustworthy_ai_2019).
-These principles, formulated prior to the adoption of the EU AI Act, have long served as a foundational reference for aligning AI development with European ethical and societal values.
-Although the AI Act introduces a more prescriptive and legally binding framework, the essence of these principles remains embedded throughout the regulation: in some cases as stand-alone articles (e.g., transparency, human oversight), and in others distributed across multiple provisions concerning governance, documentation, and system performance.
-To complement these ethical foundations with more operational and procedural dimensions, we added four additional categories corresponding to key obligations introduced by the AI Act: quality management, risk management, technical documentation, and record keeping.
-These categories address the infrastructural and organizational mechanisms that ensure traceability, consistency, and accountability across the AI lifecycle.
-Together, the original HLEG principles and these four operational domains constitute the 11 macro-requirements underpinning the proposed framework.
-Building upon this foundation, each macro-category was further refined into a set of more granular requirements.
-These were derived through a close reading of the AI Act’s legal provisions, complemented where necessary by interpretative inference from the broader normative intent of the macro-level requirements.
-This two-tiered decomposition, from principles to operational requirements, enabled a systematic and traceable linkage between high-level ethical concepts and their concrete technical or procedural manifestations.
-Once the granular requirements were defined, we proceeded to map each of them to specific guidelines and methods.
-This mapping was performed hierarchically: first connecting each requirement to relevant guidance sources, and subsequently identifying testing or control methods capable of empirically or procedurally validating compliance.
-This structured mapping approach ensures conceptual continuity across levels, from principle to requirement, from requirement to guideline, and from guideline to method, thereby providing a coherent foundation for both qualitative governance assessment and quantitative system testing.
-3.2. Sources
-The identification of control and testing methods followed a structured, multi-stage approach aimed at ensuring both regulatory relevance and technical completeness.
-The objective was to derive mechanisms that are not only aligned with current legal and ethical requirements, but also grounded in established standards and empirically validated research.
-First, we conducted a systematic examination of binding and quasi-binding regulatory sources, beginning with the EU AI Act (eu_ai_act_2024) and the General Data Protection Regulation (GDPR)  (GDPR2016).
-These texts define the normative baseline for high-risk AI systems in the European context, providing explicit legal obligations related to transparency, robustness, and accountability.
-Their provisions were used to delimit the scope of each requirement and to establish traceability between legal articles and technical mechanisms.
-Second, we reviewed international standards and authoritative guidance documents issued by globally recognised bodies, such as the EU’s General Purpose AI Code of Practice  (gpai-code-practice-3rd), ISACA’s Advanced Audit in AI Official Review Manual (isaca2025aaia), UNESCO’s Recommendation on the Ethics of Artificial Intelligence (unesco2021ethics), ISO/IEC 42001:2023 (iso_iec_42001_2023), ISO23894 (ISO23894) and ISO9001 (ISO9001).
-These documents were selected because they operationalize high-level principles into implementable governance or audit processes and therefore provide an essential bridge between normative and technical domains.
-Third, where specific mappings between legal principles and technical procedures were not sufficiently covered by regulatory or standardization sources, we extended the analysis to the scientific literature.
-We focused on highly recognised and widely cited works in AI governance, robustness testing, and risk management (e.g., (Brundage2020; Basiri2016; Floridi2018; mundhenk2020)).
-Inclusion criteria combined citation count, influence of the publication venue, and the relevance of the paper to measurable aspects of trustworthiness such as fairness, reliability, or transparency.
-This ensured that the resulting mappings draw from the most authoritative and empirically supported methodologies currently available.
-Through this layered approach, progressing from regulations to standards to scientific evidence, we aimed to guarantee both the normative validity and the technical soundness of the proposed framework.
-Each subsection that follows in Section 4 corresponds to one Trustworthy AI requirement and enumerates its associated control and testing mechanisms accordingly.
-3.3. Dimensions of Assessment
-Figure 1. Mapping between the assessment type and the assessment target dimensions.
-Before mapping requirements to methods in Section 4, it is necessary to introduce the conceptual structure underpinning the framework.
-The analysis of an AI system can be meaningfully organised along two complementary dimensions of assessment.
-The first dimension concerns the type of assessment, The second dimension concerns the target of assessment.
-Together, these two axes define the space in which assurance, verification, and risk mitigation take place.
-Along the type of assessment axis, the framework distinguishes between controls and testing.
-Controls are primarily ex-ante activities that establish governance, documentation, and procedural safeguards to ensure responsible development before deployment.
-Testing, by contrast, consists of ex-post evaluations that empirically examine system behaviour, performance, or statistical properties under specified conditions.
-Neither approach is hierarchical; rather, they are mutually reinforcing.
-Controls ensure that development processes are auditable and accountable, while testing provides empirical evidence that outcomes align with intended requirements.
-Along the target of assessment axis, the framework identifies four lifecycle layers.
-The data layer captures aspects related to data governance such as the acquisition, composition, and representativeness of data.
-Model related concerns such as its architecture, explainability, robustness, and performance are covered by the model dimension.
-The processes layer encompasses organisational and procedural elements of system development and maintenance, including risk and quality management, documentation, traceability, and lifecycle governance.
-It reflects how the system is produced and managed rather than what it does.
-Finally, the final product layer focuses on the behaviour of the deployed system, its interaction with users, and its broader societal or environmental implications, where outputs and impacts can be empirically observed and validated.
-The intersection of these two axes yields a comprehensive assessment model that connects organisational assurance with empirical verification across all lifecycle stages.
-This structure enables evaluators to determine, for each obligation under the AI Act, which combination of assessment type and target is most appropriate, ensuring that governance and verification are treated as interdependent components of a coherent assurance continuum.
+This section describes the methodology used to construct a structured mapping between high-level AI Act requirements and concrete, implementable verification activities.
+Figure 1 provides an overview of the resulting compliance verification structure and the main methodological building blocks.
+It illustrates how high-level legal requirements are progressively decomposed, grounded, and translated into concrete verification activities, which are then situated within a structured verification space defined along two core dimensions.
+Figure 1. Overview of the methodology.
+2.1. Normative inputs and requirement structuring
+The starting point of the methodology consists in identifying and structuring a set of high-level requirements derived from the obligations established by the AI Act.
+Given the heterogeneity and dispersion of these obligations across the regulation, an explicit structuring step is required to organise them into a coherent form suitable for systematic verification.
+For this purpose, we use the seven principles of Trustworthy AI defined by the European Commission’s High-Level Expert Group in the Ethics Guidelines for Trustworthy AI (eu_trustworthy_ai_2019) as an organising abstraction.
+Although these principles predate the AI Act and are not legally binding, they have strongly influenced the development of European AI governance and are reflected throughout the regulation, both as explicit obligations (e.g.
+transparency, human oversight) and as distributed requirements embedded in provisions on governance, documentation, and system performance.
+Their role in the methodology is therefore structural: they serve as a stable taxonomy for grouping and analysing legally binding requirements, not as an independent source of obligations.
+To account for the more prescriptive and organisational character of the AI Act, this abstraction is complemented with four additional requirement categories corresponding to obligations that are central to conformity assessment for high-risk systems: quality management, risk management, technical documentation, and record keeping.
+These categories capture lifecycle governance and accountability mechanisms that are explicitly mandated by the regulation but are only partially represented in principle-based formulations.
+Together, the resulting eleven high-level requirements provide a structured normative representation of the obligations imposed by the AI Act for high-risk systems.
+Each requirement corresponds to a coherent cluster of legally binding provisions, organised in a form suitable for systematic verification and without introducing additional normative content beyond the regulation itself.
+Each high-level requirement is subsequently decomposed into more granular operational components.
+These components correspond either to obligations explicitly stated in the AI Act or to requirements that can be reasonably derived from its provisions in order to enable verification.
+This decomposition does not aim to provide an exhaustive or definitive interpretation of the law, but to articulate traceable operational interpretations that can be subjected to assessment.
+The methodology explicitly allows for alternative decompositions depending on context, sector, and risk profile, and is designed to accommodate such variability rather than eliminate it.
+2.2. Authoritative grounding
+To ensure both normative alignment and technical soundness, the operational interpretation of requirements is grounded in authoritative sources that specify or exemplify how high-level obligations are implemented and assessed in practice.
+These sources are predominantly non-binding, complemented where appropriate by binding instruments whose established operational mechanisms are widely reused in compliance and assurance activities.
+In this respect, the GDPR (GDPR2016) is included not as a source of normative extension of the AI Act, but because it provides mature and widely operationalised mechanisms—such as risk-based assessment, documentation practices, and accountability procedures—that directly inform how several AI Act obligations are interpreted and verified in practice.
+The AI Act explicitly operates in complementarity with existing data protection law, and high-risk AI systems in the European context frequently involve personal data processing throughout development, deployment, and monitoring.
+As a result, verification activities related to data governance, transparency and information provision, record keeping, and affected-person rights often rely on GDPR-derived concepts, artefacts, and procedures, such as information duties, accountability documentation, and impact-assessment practices.
+For this reason, the GDPR is treated here as an authoritative operational reference within the grounding layer.
+Beyond this exception, operational grounding relies on non-binding but authoritative sources that provide reusable mechanisms for implementing and assessing regulatory obligations in practice.
+These include the EU General Purpose AI Code of Practice (gpai-code-practice-3rd), ISACA’s Advanced Audit in AI Official Review Manual (isaca2025aaia), UNESCO’s Recommendation on the Ethics of AI (unesco2021ethics), the Web Content Accessibility Guidelines (WCAG) 2.2 (wcag22), and a broad range of ISO and IEC standards relevant to AI governance, risk management, quality management, cybersecurity, and data quality (e.g.
+(ISO_IEC_22989_2022; ISO23894; ISO_IEC_TR_24027_2021; ISO_IEC_24029_2_2023; ISO_IEC_24970_DIS_2025; ISO_IEC_27001_2022; ISO_IEC_27005_2022; ISO_IEC_27035_1_2023; ISO9001; ISO_31000_2018; ISO_IEC_31010_2019; iso_iec_42001_2023; ISO_IEC_5259_1_2024; ISO_8000_8_2015)).
+These sources were selected because they translate high-level principles into implementable governance processes, controls, and audit practices, thereby bridging normative requirements and operational verification.
+Where regulatory, guidance, and standardisation sources did not sufficiently specify technical verification procedures, the analysis was extended to the scientific literature.
+Highly cited and influential works in AI governance, robustness testing, risk management, and human–system interaction were considered (e.g.
+(Brundage2020; Basiri2016; Floridi2018; mundhenk2020; sheridan1992telerobotics; schuler1993participatory; CoE_AI_2020)).
+Selection criteria combined citation impact, venue relevance, and explicit connection to measurable or testable aspects of AI trustworthiness.
+This layered grounding process ensures that each operational requirement and associated verification activity can be traced back to authoritative legal, standardisation, or scientific sources.
+2.3. Verification dimensions
+Figure 2. Verification space defined by verification type and verification target.
+Concrete verification activities are organised within a structured verification space defined by two orthogonal dimensions: the type of verification and the target of verification within the AI lifecycle.
+Together, these dimensions provide a common operational language for identifying, organising, and comparing verification activities across heterogeneous systems and stakeholders.
+The first dimension concerns the type of verification.
+Two broad categories are distinguished: controls and testing.
+Controls refer to process-based assurance mechanisms, including governance structures, documentation practices, quality management procedures, and organisational safeguards that ensure responsible system development and operation.
+Testing refers to empirical evaluations that assess system behaviour, performance, or properties under specified conditions.
+Controls and testing are complementary rather than hierarchical: controls establish accountability and traceability, while testing provides empirical evidence of compliance.
+Both may be applied at any stage of the AI lifecycle.
+The second dimension concerns the target of verification, corresponding to the lifecycle component or artefact being evaluated.
+Numerous software and AI lifecycle models provide fine-grained decompositions of development and deployment phases (e.g.
+(de2022artificial; wang2021artificial)).
+However, such models vary significantly across standards, engineering practices, and organisational contexts, and their level of detail often presupposes technical expertise.
+For the purpose of compliance verification, which involves regulators, auditors, legal experts, risk managers, and developers alike, a higher level of abstraction is required to support immediate interpretability and shared understanding.
+Accordingly, the methodology adopts four high-level verification targets that act as stable and intuitive aggregation buckets rather than exhaustive lifecycle stages.
+The data target addresses issues related to data acquisition, composition, representativeness, and quality.
+The model target concerns the AI system’s architecture, training procedures, explainability, robustness, and performance characteristics.
+The processes target encompasses organisational and procedural elements such as risk management, quality assurance, documentation, traceability, and lifecycle governance.
+The final product target focuses on the behaviour of the deployed system, including its interaction with users, integration into sociotechnical contexts, and observable outputs and impacts.
+This abstraction is intentional: it sacrifices fine-grained lifecycle specificity in favour of clarity, stability, and cross-disciplinary usability.
+Each target is sufficiently broad to accommodate different development methodologies and technical implementations, while remaining concrete enough to support systematic verification and evidence collection.
+While analytically distinct, these targets may overlap in practice.
+Certain verification activities, such as model monitoring or data versioning, span multiple targets simultaneously.
+The methodology treats these boundaries as pragmatic rather than absolute, allowing verification activities to address multiple concerns where appropriate.
+Within this verification space, each high-level requirement is associated with a set of control and testing activities consistent with its legal intent and operational scope.
+This structure enables evaluators to systematically identify relevant verification activities, organise evidence coherently, and ensure coverage across governance and empirical dimensions.
+By making explicit how regulatory requirements translate into concrete verification actions, the methodology supports comparability, traceability, and communication across heterogeneous stakeholders and use cases.
 3. Mapping
-This section presents the mapping between the macro-requirements of Trustworthy AI and the corresponding mechanisms that enable their implementation and verification.
-Each subsection (R1–R11) details one requirement category, identifying the relevant legal provisions, associated guidelines, and the methods through which compliance can be assessed or demonstrated.
-To ensure conceptual clarity, each method is classified along the two analytical dimensions described in 3.
-Type distinguishes between C (Controls) and T (Tests).
-Target specifies the primary layer of the AI lifecycle to which the method applies: D (Data), M (Models) , P (Processes) , and FP (Final Product).
+In this section, we present the mapping between the high-level requirements of the AI Act defined in Section 2.1 and the mechanisms that support their implementation and verification.
+Each subsection (R1–R11) introduces a brief explanation of the corresponding requirement, followed by the actual mapping presented in the form of a table.
+These tables identify the relevant legal provisions, associated authoritative references, and the concrete methods through which compliance can be assessed or demonstrated.
+To ensure conceptual clarity and consistency across requirements, each method is classified along the two analytical dimensions described in Section 2.3.
+Type distinguishes between C (Controls) and T (Tests), while Target specifies the primary layer of the AI lifecycle to which the method applies: D (Data), M (Models), P (Processes), and FP (Final Product).
 3.1. R1: Human Agency and Oversight
@@
 In practice, this means that individuals supervising the system must understand its capabilities and limitations, be able to intervene when necessary, and override its decisions to prevent or mitigate risks.
-The article also reinforces user autonomy by mandating transparency, informed consent, and the possibility to opt out of AI-driven functionalities.
-Together, these obligations ensure that human judgment and accountability remain central throughout the design, deployment, and use of AI systems.
-Requirement Guidelines Methods Type Target
-R1.1: Human in the loop AIA (Art. 14.4.d–e) M1.1: Manual override — Confirm that a human operator can take control of the AI system at any time during operation. C P, FP
-AIA (Art. 14.4.a–c) M1.2: Supervisory control — Assess whether human operators are able to supervise the AI system and intervene when necessary. C P, FP
+Requirement References Methods Type Target
+R1.1: Human in the loop Sheridan et al. (sheridan1992telerobotics) M1.1: Manual override — Confirm that a human operator can take control of the AI system at any time during operation. C P, FP
+Sheridan et al. (sheridan1992telerobotics) M1.2: Supervisory control — Assess whether human operators are able to supervise the AI system and intervene when necessary. C P, FP
 ISO/IEC 42001:2023 (Ann.A.3.2) M1.3: Strategic governance — Ensure that organisational oversight mechanisms for AI deployment are defined and documented. C P
-R1.2: User Autonomy AIA (Art. 50.1) M1.4: Informed consent — Check that users are provided with clear information about AI usage and are able to give or withhold consent. C FP
-ISACA manual (Sec. 3.2, p.137) M1.5: User preferences — Review whether the system allows adaptation of its behaviour based on user-set preferences. C FP
-ISACA manual (Sec. 3.2, p.137) M1.6: Opt-in/out — Confirm that users can enable or disable AI functionalities according to their choices. C FP
-R1.3: Oversight Mechanisms ISACA manual (Sec. 2.11.1) M1.7: Audit trails — Ensure that system actions and decisions are recorded in tamper-evident logs for traceability. C P
+R1.2: User Autonomy GDPR (Art.7); Andreotta et al.
+(andreotta2022ai) M1.4: Informed consent — Check that users are provided with clear information about AI usage and are able to give or withhold consent.
+C FP
+ISACA manual (§ 3.2) M1.5: User preferences — Review whether the system allows adaptation of its behaviour based on user-set preferences. C FP
+ISACA manual (§ 3.2) M1.6: Opt-in/out — Confirm that users can enable or disable AI functionalities according to their choices. C FP
+R1.3: Oversight Mechanisms ISACA manual (§ 2.11.1) M1.7: Audit trails — Ensure that system actions and decisions are recorded in tamper-evident logs for traceability. C P
 3.2. R2: Technical Robustness and Safety
@@
 They also mandate mechanisms for fallback and uncertainty management, so that any deviation from intended operation can be detected, controlled, and mitigated in a way that preserves both functionality and safety.
-Requirement Guidelines Methods Type Target
-R2.1: Resilience & Reliability ISACA manual (Sec. 1.4.4) M2.1: Stress testing — Assess whether the system has been exposed to extreme loads or edge cases to identify potential failure points. T P
-ISACA manual (Sec. 2.8.6) M2.2: Drift detection — Confirm that mechanisms are in place to monitor data distribution changes and evaluate behaviour under such shifts. T D, P
+Requirement References Methods Type Target
+R2.1: Resilience & Reliability ISACA manual (§ 1.4.4) M2.1: Stress testing — Assess whether the system has been exposed to extreme loads or edge cases to identify potential failure points. T P
+ISACA manual (§ 2.8.6) M2.2: Drift detection — Confirm that mechanisms are in place to monitor data distribution changes and evaluate behaviour under such shifts. T D, P
 Basiri et al. (Basiri2016) M2.3: Chaos engineering — Determine whether resilience has been tested through controlled fault injection in production-like environments. T P
-AIA (Art. 15.4); ISACA manual (Sec. 2.9.4) M2.4: Fail-over architecture — Ensure redundant components exist to maintain functionality in case of failure. C P, FP
-R2.2: Robustness to Attacks Brundage et al. (Brundage2020); ISACA manual (Sec. 2.13.3) M2.5: Red teaming — Examine whether adversarial testing is conducted to uncover vulnerabilities. T P, FP
-R2.3: Fallback & Fail-safe HLEG Guidelines; AIA (Art. 14.4.e) M2.6: Human fallback — Check that the system includes a mechanism to transfer control to a human in the event of failure. C P, FP
-R2.4: Accuracy & Uncertainty ISACA manual (Sec. 1.4.3) M2.7: Model calibration — Review whether confidence scores are appropriately aligned with model performance. T M
+ISACA manual (§ 2.9.4) M2.4: Fail-over architecture — Ensure redundant components exist to maintain functionality in case of failure. C P, FP
+R2.2: Robustness to Attacks Brundage et al. (Brundage2020); ISACA manual (§ 2.13.3) M2.5: Red teaming — Examine whether adversarial testing is conducted to uncover vulnerabilities. T P, FP
+R2.3: Fallback & Fail-safe HLEG Guidelines M2.6: Human fallback — Check that the system includes a mechanism to transfer control to a human in the event of failure. C P, FP
+R2.4: Accuracy & Uncertainty ISACA manual (§ 1.4.3) M2.7: Model calibration — Review whether confidence scores are appropriately aligned with model performance. T M
 3.3. R3: Privacy and Data Governance
 Article 10 of the AI Act establishes specific obligations concerning data governance and data quality for high-risk AI systems, requiring that datasets be relevant, representative, free of errors, and complete to the extent possible.
-These provisions complement the data protection principles set out in the GDPR, particularly Articles 5, 25, and 32, which enshrine purpose limitation, privacy by design, and security of processing.
-Together, they ensure that data is collected, processed, and stored in a lawful, fair, and transparent manner.
-The combined framework mandates both organisational and technical measures—such as encryption, pseudonymisation, and role-based access control—to prevent unauthorised use and ensure accountability for data-related decisions.
-Effective data governance thus underpins not only compliance with legal obligations but also the technical robustness and trustworthiness of AI systems across their lifecycle.
-Requirement Guidelines Methods Type Target
+These obligations are AI-specific and apply irrespective of whether personal data is involved.
+Where high-risk AI systems rely on personal data, the operational interpretation and verification of Article 10 requirements necessarily draw on established data protection mechanisms defined under the GDPR, in particular Articles 5, 25, and 32, which address purpose limitation, data protection by design, and security of processing.
+In such cases, GDPR concepts and safeguards provide concrete, widely operationalised mechanisms through which certain data governance obligations of the AI Act can be implemented and assessed in practice.
+Accordingly, compliance with Article 10 typically involves a combination of organisational and technical measures, such as encryption, pseudonymisation, access controls, and accountability procedures, which are already well established in data protection practice.
+Effective data governance thus supports compliance with AI Act requirements while also contributing to the technical robustness and trustworthiness of AI systems across their lifecycle.
+Requirement References Methods Type Target
 R3.1: Data Minimisation GDPR (Art. 5.1.b–c) M3.1: Purpose limitation — Verify that only data strictly necessary for a defined processing purpose is collected. C D, P
@@
 R3.3: Access & Security GDPR (Art. 32.1.b–d) M3.4: Role-based access control — Ensure access rights are granted based on organisational roles. C P
-GDPR (Art. 32.1.a); ISACA manual (Sec. 2.7.3) M3.5: Encryption — Check that data is encrypted in storage and transmission. T D, P
-R3.4: Data Provenance ISACA manual (Sec. 2.7.5) M3.6: Lineage tracking — Confirm that data transformations are logged to ensure traceability. C D, P
-ISACA manual (Sec. 3.1.2) M3.7: Metadata versioning — Verify dataset versions and schema changes are tracked. C D, P
+GDPR (Art. 32.1.a); ISACA manual (§ 2.7.3) M3.5: Encryption — Check that data is encrypted in storage and transmission. T D, P
+R3.4: Data Provenance ISACA manual (§ 2.7.5) M3.6: Lineage tracking — Confirm that data transformations are logged to ensure traceability. C D, P
+ISACA manual (§ 3.1.2) M3.7: Metadata versioning — Verify dataset versions and schema changes are tracked. C D, P
 R3.5: Consent Management GDPR (Art. 7.1–2) M3.8: Consent tracking — Inspect whether user consents are explicitly obtained, logged, and tied to specific purposes. C P, FP
@@
 Transparency is one of the most fundamental requirements for trustworthy and accountable AI.
-Under the AI Act, AI systems must be designed and documented in a way that allows users, regulators, and auditors to understand their capabilities, limitations, and decision-making logic.
+Under the AI Act, specifically in Art.
+13, AI systems must be designed and documented in a way that allows users, regulators, and auditors to understand their capabilities, limitations, and decision-making logic.
 Transparency encompasses both technical and communicative aspects: it includes model explainability, data and process traceability, disclosure of AI usage to end users, and interpretability of outputs in accessible language.
 Together, these mechanisms ensure that AI operations remain observable and interpretable across all stages of their lifecycle, thereby fostering accountability and enabling meaningful human oversight.
-Requirement Guidelines Methods Type Target
-R4.1: Explainability ISACA manual (Sec.
-3.2, p.139) M4.1: Local interpretability — Verify that the system provides human-understandable explanations of model outputs using recognised local explanation techniques (e.g., feature attribution or surrogate models).
+Requirement References Methods Type Target
+R4.1: Explainability ISACA manual (§ 3.2) M4.1: Local interpretability — Verify that the system provides human-understandable explanations of model outputs using recognised local explanation techniques (e.g., feature attribution or surrogate models).
 T D, M, FP
-Mundhenk et al. (mundhenk2020) M4.2: Saliency maps — Check if visualization tools highlight relevant input regions. T D, M, FP
-R4.2: Model & Data Traceability GPAI CoP (Sec. Transparency) M4.3: Model cards — Ensure that standardized documentation is available and maintained. C M, P
+Mundhenk et al. (mundhenk2020) M4.2: Saliency maps — Check if visualisation tools highlight relevant input regions. T D, M, FP
+R4.2: Model & Data Traceability GPAI CoP (Sec. Transparency) M4.3: Model cards — Ensure that standardised documentation is available and maintained. C M, P
 GPAI CoP (Sec. Transparency) M4.4: Datasheets — Assess whether datasheets describing dataset origin and composition are accessible. C D, P
-R4.3: Disclosure of AI Use AIA (Art. 50.1) M4.5: AI identity notices — Verify that users are clearly informed when interacting with an AI system. C FP
-GPAI CoP; ISACA manual (Sec. 3.2, p.139) M4.6: Disclaimers — Check whether limitations and assumptions are transparently disclosed. C FP
-AIA (Art. 13.3.b.III) M4.7: Risk communication — Assess whether the system communicates known risks and uncertainties. C FP
-R4.4: Interpretability for End Users AIA (Art. 13.2) M4.8: Understandable language — Confirm that technical terms are explained using plain language. C FP
+R4.3: Disclosure of AI Use ISO/IEC 22989 (§5.15) M4.5: AI identity notices — Verify that users are clearly informed when interacting with an AI system. C FP
+GPAI CoP; ISACA manual (§ 3.2) M4.6: Disclaimers — Check whether limitations and assumptions are transparently disclosed. C FP
+ISO/IEC 23894 (§6.1) M4.7: Risk communication — Assess whether the system communicates known risks and uncertainties. C FP
+R4.4: Interpretability for End Users ISO/IEC 22989 (§5.15) M4.8: Understandable language — Confirm that technical terms are explained using plain language. C FP
 3.5. R5: Diversity, Non-discrimination and Fairness
 Diversity, non-discrimination, and fairness are central to ensuring that AI systems respect fundamental rights and avoid reinforcing structural biases.
-Articles 9, 10, and 27 of the AI Act require that systems be designed, trained, and deployed in a way that prevents discriminatory outcomes and ensures inclusive access.
+Articles 10, and 27 of the AI Act require that systems be designed, trained, and deployed in a way that prevents discriminatory outcomes and ensures inclusive access.
 This includes assessing data representativeness, applying fairness metrics across protected groups, and incorporating participatory design practices that reflect societal diversity.
 Fairness thus spans both technical and organisational domains, linking data quality, model evaluation, and stakeholder engagement to the ethical and legal accountability of AI systems.
-Requirement Guidelines Methods Type Target
-R5.1: Fairness in Data ISACA manual (Sec. 3.2, p.138) M5.1: Bias detection — Evaluate whether datasets are examined for representation gaps. T D, P
-ISACA manual (Sec. 3.2, p.138) M5.2: Sampling strategies — Confirm that sampling ensures balanced representation. C D, P
-R5.2: Fairness in Algorithms ISACA manual (Sec. 3.2, p.138) M5.3: Protected-class fairness metrics — Verify that fairness metrics are applied across protected groups. T D, M, P
-R5.3: Inclusive Design AIA (Rec. 132) M5.4: Accessibility — Check that the system supports assistive technologies. C FP
-R5.4: Stakeholder Engagement AIA (Rec. 165) M5.5: Participatory design — Assess whether affected users and communities are involved in design and testing. C FP, P
+Requirement References Methods Type Target
+R5.1: Fairness in Data ISACA manual (§ 3.2) M5.1: Bias detection — Evaluate whether datasets are examined for representation gaps. T D, P
+ISACA manual (§ 3.2) M5.2: Sampling strategies — Confirm that sampling ensures balanced representation. C D, P
+R5.2: Fairness in Algorithms ISACA manual (§ 3.2) M5.3: Protected-class fairness metrics — Verify that fairness metrics are applied across protected groups. T D, M, P
+R5.3: Inclusive Design WCAG 2.2 M5.4: Accessibility — Check that the system supports assistive technologies. C FP
+R5.4: Stakeholder Engagement Schuler et al. (schuler1993participatory) M5.5: Participatory design — Assess whether affected users and communities are involved in design and testing. C FP, P
 3.6. R6: Societal and Environmental Well-being
@@
 These requirements embed sustainability and societal benefit as integral dimensions of AI governance.
-Requirement Guidelines Methods Type Target
-R6.1: Environmental Impact UNESCO (Rec. 84); ISACA manual (Sec. 1.17.7) M6.1: Energy-use tracking — Verify that the system monitors computational energy consumption and carbon footprint. T P, FP
-R6.2: Ethical Alignment AIA (Art. 27.1) M6.2: Rights-based system — Assess whether system design aligns with fundamental rights. C P
+Requirement References Methods Type Target
+R6.1: Environmental Impact UNESCO (Rec. 84); ISACA manual (§ 1.17.7) M6.1: Energy-use tracking — Verify that the system monitors computational energy consumption and carbon footprint. T P, FP
+R6.2: Ethical Alignment Leslie et al. (CoE_AI_2020) M6.2: Rights-based system — Assess whether system design aligns with fundamental rights. C P
 UNESCO (Rec. 58) M6.3: Independent ethics committee — Verify that an independent ethics committee or officer is appointed. C P
@@
 They require clear responsibility allocation, secure logging, version-controlled documentation, and channels for incident reporting, thereby operationalising the principle that accountability cannot be delegated to the machine.
-Requirement Guidelines Methods Type Target
-R7.1: Responsibility Assignment ISACA manual (Sec. 3.2, p.129) M7.1: RASCI charts — Confirm that responsibility matrices define roles for each lifecycle activity. C P
-R7.2: Auditability ISO/IEC 42001:2023 (Ann.B.6.2.8); ISACA manual (Sec. 2.11.1) M7.2: Logging — Verify that detailed, tamper-evident logs of system operations are maintained and reviewed. C P
+Requirement References Methods Type Target
+R7.1: Responsibility Assignment ISACA manual (§ 3.2) M7.1: RASCI charts — Confirm that responsibility matrices define roles for each lifecycle activity. C P
+R7.2: Auditability ISO/IEC 42001:2023 (Ann.B.6.2.8); ISACA manual (§ 2.11.1) M7.2: Logging — Verify that detailed, tamper-evident logs of system operations are maintained and reviewed. C P
 ISO/IEC 42001:2023 (Ann.B.8.3) M7.3: External reporting interfaces — Ensure that secure portals are available for third-party auditors. C P
 R7.3: Incident Reporting Floridi et al. (Floridi2018) M7.4: Redress systems — Check whether users can report issues or contest AI-driven outcomes. C FP, P
-R7.4: Documentation Integrity ISO/IEC 42001:2023 (Ann.B.6.2.5); ISACA manual (Sec. 2.11.1) M7.5: Version-controlled records — Assess whether documentation is versioned and traceable over time. C P
+R7.4: Documentation Integrity ISO/IEC 42001:2023 (Ann.B.6.2.5); ISACA manual (§ 2.11.1) M7.5: Version-controlled records — Assess whether documentation is versioned and traceable over time. C P
 3.8. R8: Quality Management
@@
 The QMS ensures that the organisation applies consistent procedures for design, testing, data management, and post-market monitoring.
-It provides a structured approach for achieving and maintaining conformity with the requirements of the AI Act, aligning with recognised standards such as ISO 9001 and ISO/IEC 42001.
+It provides a structured approach for achieving and maintaining conformity with the requirements of the AI Act.
 Quality management thereby supports traceability, continuous improvement, and long-term compliance across the AI system lifecycle.
-Requirement Guidelines Methods Type Target
+Requirement References Methods Type Target
 T8.1: Quality policy & objectives ISO 9001:2015 M8.1: Quality Policy Declaration - Verify that top management has issued a formal, documented policy defining quality objectives and regulatory compliance commitments for high-risk AI systems under a Quality Management System (QMS).
@@
 It must ensure that residual risks are acceptable relative to the intended use and that appropriate mitigation measures are applied.
-The methodology used here aligns with a set of horizontal, cross-sectoral standards, including ISO/IEC 23894 (AI risk management), ISO/IEC 42001 (AI management system), ISO 31000 and ISO 31010 (General risk management principles and techniques), as well as data- and robustness-related standards such as ISO/IEC 5259, ISO 8000, ISO/IEC TR 24027, ISO/IEC 24029, and ISO/IEC 24970.
-In coming days, when harmonized standards under the AI Act, particularly prEN 18228 (AI Risk Management) and prEN 18286 (AI Quality Management System) will be finalized, they will provide the solid structure for conformity assessment and presumption of conformity once a system is categorized as high-risk.
-However, it is essential to underline that none of these standards, both international ISO/IEC standards and future harmonized European standards, determine whether an AI system is classified as high-risk under Article 6.
-Classification is governed exclusively by the legal criteria of the AI Act, particularly Annex I and Annex III.
+The methodology used here aligns with a set of horizontal, cross-sectoral standards, including ISO/IEC 23894 (AI risk management), ISO/IEC 42001 (AI management system), ISO 31000 and ISO 31010 (General risk management principles and techniques), ISO/IEC 22989 (AI concepts and terminology) as well as data- and robustness-related standards such as ISO/IEC 5259 series, ISO 8000 series, ISO/IEC TR 24027, ISO/IEC 24029-1/-2, and ISO/IEC 24970.
+In the future, when harmonised standards under the AI Act, particularly prEN 18228 (AI Risk Management) and prEN 18286 (AI Quality Management System) will be finalised, they will provide the solid structure for conformity verification and presumption of conformity once a system is categorised as high-risk.
+However, it is essential to underline that none of these standards, both international ISO/IEC standards and future harmonised European standards, determine whether an AI system is classified as high-risk under Article 6.
+Classification is governed exclusively by the legal criteria of the AI Act, particularly in Annex III.
 Standards are used in this report solely to support structured analysis, documentation and evidence generation, and they become fully applicable only after an AI system has been classified as high-risk.
@@
 These frameworks are highly relevant within their respective regulatory regimes but are not part of the horizontal, cross-sector risk-classification process defined by the AI Act and therefore are not included in the present analysis.
-Requirement Guidelines Methods Type Target
-T9.1: System Definition ISO/IEC 42001 (§§ 4.3, 6.1.2); ISO/IEC 23894 § 5.3 M9.1 Description of AI system – define intended purpose for Art 6 mapping and record system details C P
-T9.2: Risk identification ISO/IEC 42001 (§§ 6.1.2, 7.5); ISO/IEC 23894 § 5.4; ISO 31000; ISO 31010; ISO/IEC 5259; ISO 8000; ISO/IEC TR 24027 M9.2 Structured risk identification feeding both Art.
-6 classification and Art.
-9 risk management C P
-T9.3: Risk analysis ISO/IEC 23894; ISO/IEC 42001; ISO 31000; ISO 31010; M9.3 Core classification step – apply AI Act Art. 6: Annex I & Annex III gates based on structured analysis C P
-T9.4: Risk evaluation & prioritisation ISO/IEC 23894; ISO/IEC 42001; ISO 31000; ISO 31010 M9.4 Finalise classification decision and justification recording C P
-T9.5: Risk control measures ISO/IEC 23894 §§ 5.5–5.7; ISO/IEC 42001; ISO 31000; ISO/IEC 24029-1/-2; ISO/IEC TR 24027; ISO/IEC 5259; ISO 8000; ISO/IEC 27001; 27005 M9.5 Risk control strategy documented and aligned with AI Act essential requirements C FP
-T9.6: PMM: Planning ISO/IEC 42001 (§§ 8.4, 9.1, 7.5); ISO/IEC 23894 (§§ 5.3, 5.7); ISO 31000; ISO/IEC 24970; ISO/IEC 27001; 27005 M9.6 PMM system established with structured monitoring requirements C P
-T9.7: PMM: Data collection & analysis ISO/IEC 42001 (§§ 9.1, 9.3, 7.5); ISO/IEC 23894 (§§ 5.5–5.7); ISO 31000; ISO/IEC 24970; ISO/IEC 27001; 27005 M9.7 Continuous assessment and early detection of anomalies or new hazards T FP
-T9.8: PMM: Review & CAPA ISO/IEC 42001 (§§ 8.3, 10, 7.5); ISO/IEC 23894 (§§ 5.6, 5.8); ISO 31000 M9.8 CAPA implemented, system updated, continuous compliance maintained C P
-T9.9: Incident reporting ISO/IEC 42001 (§§ 8.4, 9.1, 7.5); ISO/IEC 23894 (§§ 5.7, 5.8); ISO/IEC 24970; ISO/IEC 27035 M9.9 Compliant incident reporting + full documentation + integration of learnings C P
+Requirement References Methods Type Target
+T9.1: System Definition ISO/IEC 42001 (§§ 4.3, 6.1.2); ISO/IEC 23894 (§ 6.3); ISO/IEC 22989 M9.1 Description of AI system – define intended purpose for Art 6 mapping and record system details C P
+T9.2: Risk identification ISO/IEC 42001 (§§ 6.1.2, 7.5); ISO/IEC 23894 (§§ 5.4, 6.4); ISO 31000; ISO 31010; ISO/IEC 5259-1; ISO 8000-8; ISO/IEC TR 24027 M9.2 Structured risk identification relevant to classification (Art.
+6) and risk management (Art.
+9) C P
+T9.3: Risk analysis ISO/IEC 23894 (§§ 6.4, 6.5); ISO/IEC 42001; ISO 31000; ISO 31010; ISO/IEC 24029-2 M9.3 Core classification step – apply AI Act Art.
+6: Annex I & Annex III gates using structured analysis C P
+T9.4: Risk evaluation & prioritisation ISO/IEC 23894 (§§ 6.4, 6.5); ISO/IEC 42001; ISO 31000; ISO 31010 M9.4 Finalize classification decision;record justification and prioritise risks C P
+T9.5: Risk control considerations ISO/IEC 23894 (§§ 5.5–5.7); ISO/IEC 42001; ISO 31000; ISO/IEC 24029-1/-2; ISO/IEC TR 24027; ISO/IEC 5259; ISO 8000; ISO/IEC 27001; ISO/IEC 27005 M9.5 Outline potential risk areas aligned with AI Act essential requirements C FP
+T9.6: PMM: Planning ISO/IEC 42001 (§§ 8.4, 9.1, 7.5); ISO/IEC 23894 (§§ 5.3, 5.7); ISO 31000; ISO/IEC 24970; ISO/IEC 27001; ISO/IEC 27005 M9.6 Establishing PMM plan with monitoring, logging and reporting C P
+T9.7: PMM: Data collection & analysis ISO/IEC 42001 (§§ 9.1, 9.3, 7.5); ISO/IEC 23894 (§§ 5.5–5.7); ISO 31000; ISO/IEC 24970; ISO/IEC 27001; ISO/IEC 27005 M9.7 Continuous verification and early detection of anomalies or new hazards T FP
+T9.8: PMM: Review & CAPA ISO/IEC 42001 (§§ 8.3, 10, 7.5); ISO/IEC 23894 (§§ 5.6, 5.8); ISO 31000 M9.8 Perform CAPA, update system, maintain continuous compliance C P
+T9.9: Incident reporting ISO/IEC 42001 (§§ 8.4, 9.1, 7.5); ISO/IEC 23894 (§§ 5.7, 5.8); ISO/IEC 24970; ISO/IEC 27035 M9.9 Compliant incident reporting; documentation; integrate lessons learned C P
 3.10. R10: Technical Documentation
@@
 This documentation must provide sufficient detail to assess the system’s conformity, including its intended purpose, design specifications, data sources, testing methods, and risk management processes.
-Proper documentation ensures transparency, enables traceability of design choices, and supports both internal governance and external conformity assessment.
-Requirement Guidelines Methods Type Target
-R10.1: System overview & intended purpose prEN 18286; ISO 9001; ISO/IEC 42001 M10.1 System overview and intended purpose documented in alignment with Annex IV and QMS requirements. C FP
-R10.2: Design specifications & architecture prEN 18286; ISO 9001; ISO/IEC 42001 M10.2 Architecture and design specifications captured with complete traceability to risks and requirements. C M, P
-R10.3: Dataset description & provenance prEN 18286; ISO 9001; ISO/IEC 42001 M10.3 Dataset provenance, quality attributes and limitations documented in compliance with Art. 10 and Annex IV. C D
-R10.4: Performance metrics & validation prEN 18286; ISO 9001; ISO/IEC 42001 M10.4 Performance, robustness and validation evidence compiled and traceable to requirements, risks and intended purpose.
+Proper documentation ensures transparency, enables traceability of design choices, and supports both internal governance and external conformity verification.
+Requirement References Methods Type Target
+R10.1: System overview & intended purpose ISO/IEC 42001 (§§ 4.3, 6.1.2, 7.5); ISO/IEC 23894 (§  6.3); ISO/IEC 22989 M10.1 System overview and intended purpose documented in alignment with Annex IV and QMS requirements.
+C FP
+R10.2: Design specifications & architecture ISO/IEC 42001(§§  8.1–8.3, 7.5); ISO/IEC 23894 (§§  6.4–6.5) M10.2 Architecture and design specifications captured with complete traceability to risks and requirements.
+C D&M
+R10.3: Dataset description & provenance ISO/IEC 42001 (§§  7.5, 8.3); ISO/IEC 5259-1; ISO 8000-8; ISO/IEC TR 24027 M10.3 Dataset provenance, quality attributes and limitations documented in compliance with Art.
+10–11 and Annex IV.
+C D&M
+R10.4: Performance metrics & validation ISO/IEC 42001 (§§  8.3, 9.1, 7.5); ISO/IEC 23894 (§§   6.4–6.5); ISO/IEC 24029-1/-2; ISO/IEC 24970; ISO/IEC 27001; ISO/IEC 27005 M10.4 Performance, robustness and validation evidence compiled and traceable to requirements, risks and intended purpose.
 T FP
-R10.5: Compliance evidence mapping prEN 18286; ISO 9001; ISO/IEC 42001 M10.5 Structured compliance mapping demonstrating how Annex IV and Article 17 obligations are satisfied. C P
+R10.5: Compliance evidence mapping ISO/IEC 42001 (§§  6.1, 8.1––8.4, 9.1–9.3, 10); ISO/IEC 23894 (§§ 5.3–5.8); ISO/IEC 27001; ISO/IEC 27005 M10.5 Structured compliance mapping demonstrating how Annex IV and Article 17 obligations are satisfied.
+C P
 3.11. R11: Record-keeping
 Article 18 of the AI Act mandates that providers retain automatically generated logs and relevant records to ensure traceability and accountability throughout the AI lifecycle.
-Record-keeping supports post-market monitoring, facilitates incident investigation, and provides evidence for conformity assessments.
-It must include documentation of model versions, dataset lineage, and data retention or deletion policies, consistent with organisational governance and data protection requirements.
-Requirement Guidelines Methods Type Target
-R11.1: Operational logs prEN 18286; ISO 9001; ISO/IEC 42001 M11.1 Operational logging system in place with records supporting Art. 12 transparency and PMM requirements. C FP
-R11.2: Version control records prEN 18286; ISO 9001; ISO/IEC 42001 M11.2 Version control ensures full traceability and reproducibility of AI system states over time. C M
-R11.3: Model & dataset lineage prEN 18286; ISO 9001; ISO/IEC 42001 M11.3 Full lineage enables auditing, incident analysis, risk traceability and evidence-based updates. C D, M
-R11.4: Retention & deletion policies prEN 18286; ISO 9001; ISO/IEC 42001 M11.4 Retention/deletion policies implemented to ensure compliance, minimise risk and maintain only necessary records. C P
-5. Example application
-This section illustrates how the proposed framework can be applied in practice through a concrete example drawn from a real, ongoing high-risk AI use case currently being addressed in collaboration with a Swedish automotive manufacturer.
-While the framework itself is generic and technology-agnostic, this example demonstrates how it can be used to structure assurance activities across the full lifecycle of an AI system, from design and testing to validation and documentation.
-The use case concerns an AI-based system for detecting cyberattacks within connected vehicles.
-Modern vehicles rely on internal communication networks that enable electronic components to exchange information in real time.
-Protecting these networks against malicious interference is critical, as cyberattacks may compromise both vehicle safety and reliability.
-To mitigate these risks, automotive manufacturers increasingly deploy Intrusion Detection Systems that monitor in-vehicle communications and identify abnormal or potentially harmful behaviour.
-In this setting, the IDS employs AI techniques to analyse network traffic within the vehicle and detect patterns associated with known or emerging attacks.
-The system processes data from in-vehicle networks, extracts relevant features, and uses machine-learning models to distinguish normal operation from suspicious activity.
-To support transparency and engineering oversight, the system also integrates explainability mechanisms that allow engineers to understand why specific activities are flagged as intrusions.
-Designed for real-time deployment in vehicle control units or central gateways, this real-world use case illustrates how the proposed framework can be applied in a complex, safety-critical automotive context.
-In particular, the framework enables a structured and traceable assessment of data handling, model behaviour, explainability, and documentation, providing evidence to support safety, cybersecurity, and regulatory requirements.
-Legal Requirement (AI Act) Test / Control Type Method Instantiation in the Use Case
+Record-keeping supports post-market monitoring, facilitates incident investigation, and provides evidence for conformity verifications.
+It must include documentation of model versions, dataset lineage, and data retention or deletion policies, consistent with organisational governance and data protection requirements, including GDPR.
+Requirement References Methods Type Target
+R11.1: Operational logs ISO/IEC 42001 (§§  7.5, 8.3, 9.1); ISO/IEC 24970 M11.1 Operational logging system in place with records supporting Art. 18 record-keeping and PMM requirements. C FP
+R11.2: Version control records ISO/IEC 42001 (§§  7.5, 8.3) M11.2 Version control ensures full traceability and reproducibility of AI system states over time. C D&M
+R11.3: Model & dataset lineage ISO/IEC 42001 (§§  7.5, 8.3); ISO/IEC 23894 (§§ 5.3–5.8) M11.3 Full lineage enables auditing, incident analysis, risk traceability and evidence-based updates. C D&M
+R11.4: Retention & deletion policies ISO/IEC 42001 (§§  7.5, 8.3); ISO/IEC 23894 (§§ 5.3–5.8); ISO/IEC 27001; ISO/IEC 27005; GDPR (Art.
+5.1.c, e) M11.4 Retention/deletion policies implemented to ensure compliance, minimise risk and maintain only necessary records, in line with GDPR data minimisation and storage limitation principles C P
+4. Case Study
+This section illustrates how the mapping proposed in Section 3 can be applied in practice through a real, ongoing high-risk AI use case currently being analysed and tested in collaboration with Scania, a Swedish automotive manufacturer.
+While the mapping is generic and technology-agnostic, this example shows how it can structure assurance activities across the AI lifecycle, from design-time verification to validation and documentation.
+Table LABEL:tab:use_case presents the results of this mapping process; for the sake of brevity, it is reported only partially and serves as an illustrative example.
+The use case concerns an AI-based system for detecting cyberattacks within connected vehicles in an industrial setting at Scania.
+Modern vehicles rely on internal communication networks that enable real-time data exchange between electronic components.
+Protecting these networks is critical, as cyberattacks may compromise both vehicle safety and system reliability.
+To address these risks, automotive manufacturers increasingly deploy Intrusion Detection Systems (IDS) that monitor in-vehicle communications and identify abnormal or potentially harmful behaviour.
+In this setting, the IDS employs AI techniques to analyse in-vehicle network traffic and detect patterns associated with known or emerging attacks.
+The system processes network data, extracts relevant features, and uses machine-learning models to distinguish normal operation from suspicious activity.
+To support transparency and engineering oversight, explainability mechanisms are integrated to allow engineers to understand why specific activities are flagged as intrusions.
+Designed for real-time deployment in vehicle control units or central gateways, this use case demonstrates the application of the proposed mapping in a complex, safety-critical automotive context.
+The mapping enables structured and traceable verification of data handling, model behaviour, explainability, and documentation, providing evidence to support safety, cybersecurity, and regulatory requirements without intervening in system development.
+Req. Sub-req. Method Instantiation in the Use Case
 R1: Human Agency & Oversight (Art. 14) T1.1 Human-in-the-loop M1.1 Manual override Engineers and operators can override IDS decisions and reclassify labels during development.
@@
 T11.4 Retention & deletion policies M11.4 Retention/deletion Define retention periods for logs, dataset snapshots, and models; ensure GDPR compliance when identifiable metadata appears; enforce secure deletion mechanisms.
-6. Discussion and Conclusion
-The framework presented in Section 3 operationalises compliance by establishing explicit correspondences between legal requirements, recognised standards, and assessment protocols.
-Its stratified architecture—distinguishing controls and testing procedures across data, model, process, and product dimensions—articulates how different modes of assurance interrelate throughout the AI system lifecycle.
-This structure facilitates systematic self-assessment by delineating which technical and organisational elements require documented evidence of conformity, thereby supporting the structuring of documentation and monitoring activities.
-Such guidance may prove particularly valuable for organisations lacking established compliance infrastructure.
-Within regulatory sandboxes, the framework serves as a common reference architecture for aligning testing protocols and reporting practices with AI Act requirements.
-By translating legal principles into empirically verifiable criteria, it enables more transparent and structured dialogue between developers and regulatory authorities.
-Furthermore, it promotes comparability across sandbox experiments and supports coherent assessment of compliance readiness at different stages of system development.
-For formal conformity assessment, the mapping provides auditors and notified bodies with an instrument for distinguishing procedural from empirical evidence, enhancing traceability, and complementing sector-specific standards through a unified vocabulary for evaluating regulatory conformity.
-Beyond immediate applications, the framework contributes to the longer-term objective of partially or fully automated compliance verification.
-By formalising relationships between legal requirements, control mechanisms, and testing procedures, it establishes a foundation for computational systems capable of selecting, executing, and interpreting appropriate assessment protocols.
-This aligns with emerging initiatives that translate regulatory obligations into executable testing workflows and interactive monitoring dashboards (buscemi2025sandbox; ki_prueftoolsuite_2025).
-A natural extension involves developing an ontology that links each legal provision to its associated controls and testable mechanisms, thereby supporting partial automation and interoperability across assessment tools and regulatory databases.
-Incorporating a temporal dimension would additionally enable longitudinal analysis of how regulatory frameworks, assessment methodologies, and compliance expectations evolve, facilitating both retrospective evaluation and prospective governance planning.
-The framework possesses inherent limitations.
-Any mapping from legal provisions to technical methods necessarily involves interpretive judgement, and the appropriateness of individual assessment procedures remains context-dependent.
-The deliberately generic level of abstraction may constrain direct application without domain-specific adaptation.
-Consequently, the framework should be understood not as a prescriptive checklist but as a reference architecture subject to iterative refinement as standards mature and regulatory practice develops.
-Future research directions include developing prioritisation schemes among requirements, integrating continuous monitoring mechanisms, and extending the framework to encompass additional EU instruments relevant to AI governance.
-This study provides an initial architectural foundation for decomposing AI Act obligations into structured, empirically testable components.
-By establishing coherent mappings among requirements, assessment methods, and implementation guidelines, it seeks to bridge epistemic and operational divides among regulators, risk managers, developers, technical auditors, and prospective certification bodies.
-The framework is conceived as an evolving infrastructure—supporting transparent, consistent, and traceable compliance verification while fostering evidence-based dialogue between technical and legal communities.
-It represents both a methodological contribution and an invitation to collaborative development toward a shared, adaptive, and continuously improving ecosystem for trustworthy AI assurance.
+5. Discussion
+Rather than functioning solely as an analytical construct, the proposed mapping is intended to support the institutional ecosystem through which AI Act compliance is operationalised in practice.
+By linking legal requirements to identifiable and verifiable assessment activities, the mapping provides a common reference structure that can be reused across provider self-assessment, regulatory sandbox experimentation, and formal conformity assessment.
+This addresses a central implementation risk of the AI Act: that identical legal obligations may be translated into materially different verification practices across organisations, sectors, and Member States.
+For providers, the mapping enables structured self-assessment as part of continuous monitoring and post-market obligations.
+By making explicit which verification activities correspond to which legal requirements, it supports systematic internal evaluations over time, including reassessment following model updates, data drift, or changes in deployment context.
+This reduces reliance on implicit or ad hoc interpretations of compliance and facilitates the generation of traceable evidence that can be reused across internal governance, regulatory reporting, and external review.
+In regulatory sandboxes established under Article 57, the mapping can function as a shared interpretive layer between developers and competent authorities.
+By clarifying how experimental testing activities relate to specific AI Act obligations, it supports the structuring of testing plans, documentation practices, and exit reports, while preserving the flexibility required for innovation-oriented assessment.
+This role is particularly relevant in a cross-border context, where divergent sandbox practices risk emerging in the absence of a common operational reference.
+European-funded initiatives such as AI Factories (eu_ai_factories) and Testing and Experimentation Facilities (TEFs) (eu_tefs) are expected to play an increasingly central role in supporting the testing and validation of AI systems, including in the context of regulatory sandboxes, as reflected in recent policy and implementation guidance (implementingAct).
+In this setting, the availability of a shared mapping between legal requirements and verification activities becomes critical.
+These infrastructures are not merely technical providers, but intermediaries that translate regulatory expectations into concrete testing capabilities.
+A common mapping helps ensure that testing performed within such initiatives is interpretable, comparable, and reusable across regulatory and institutional contexts, rather than being tied to local or project-specific interpretations.
+More broadly, the mapping contributes to convergence in verification methodologies across the EU by offering a reusable, technology-agnostic operational vocabulary.
+While enforcement remains decentralised and sectoral specificities must be respected, convergence on underlying verification logic is essential to avoid fragmentation of compliance practices.
+The mapping is therefore intentionally designed to be extensible: it can be refined, extended, and instantiated by different actors, including competent authorities, notified bodies, EU-funded infrastructures, and sectoral initiatives.
+Such extensions can introduce domain-specific metrics or procedures while preserving traceability to the original legal requirements.
+In this sense, the mapping is best understood as an intermediate coordination layer between legal text and assessment practice.
+It does not prescribe how verification must be performed, but provides a common structure within which diverse practices can evolve in a coherent and interoperable manner.
+Encouraging its extension and sectoral instantiation is therefore not a limitation, but a necessary condition for achieving both regulatory consistency and practical relevance under the AI Act.
+Finally, while this work is grounded in the European regulatory context and explicitly targets the implementation of the AI Act, the underlying challenge it addresses is not unique to Europe.
+Many jurisdictions are currently developing or refining AI-specific regulatory frameworks that similarly rely on high-level, principle-driven obligations whose practical verification remains underspecified.
+In this respect, the approach adopted in this paper is intended to be transferable.
+The notion of an explicit mapping between legal requirements and verifiable assessment activities can be adapted to other regulatory settings, supporting regulatory learning, comparability, and institutional capacity-building beyond the EU.
+The authors therefore encourage the development of analogous mappings in other jurisdictions and see value in future cross-jurisdictional dialogue on verification methodologies as AI regulation continues to evolve globally.
+6. Conclusion
+This paper introduced an explicit and structured mapping between high-level obligations under the EU AI Act and concrete, verifiable assessment activities.
+Starting from 11 high-level requirements used as a normative basis, the mapping decomposes these obligations into 48 operational sub-requirements and associates them with 66 verification activities.
+By making these relationships explicit, the mapping bridges the gap between regulatory intent and assessable technical and organisational evidence across the AI lifecycle.
+A central contribution of this work is not merely the enumeration of verification activities, but the articulation of a reusable mapping logic.
+By organising verification activities along two orthogonal dimensions, i.e.
+the type of verification performed and the lifecycle target to which it applies, the mapping makes visible how compliance with the AI Act is constructed in practice.
+It shows that conformity emerges from structured combinations of procedural controls and empirical testing, rather than from isolated checks or single metrics.
+By formalising these relationships, the mapping directly reduces key sources of uncertainty in AI Act implementation.
+Interpretive uncertainty is addressed by clarifying how abstract legal obligations can be decomposed into operational elements grounded in recognised practices.
+Operational uncertainty is reduced by identifying verification activities that are implementable, repeatable, and auditable across different organisational and technical contexts.
+Procedural uncertainty is mitigated by enabling consistent documentation, traceability, and comparison of evidence across systems, actors, and regulatory settings.
+The value of the mapping extends beyond individual assessments.
+It provides a common reference that can be reused across provider self-assessment, regulatory sandboxes, conformity assessment, and post-market monitoring, supporting convergence in verification practices across Member States while preserving contextual flexibility.
+In doing so, it aligns with the AI Act’s governance model, which relies on continuous oversight and regulatory learning rather than one-off certification.
+The mapping is intentionally technology-agnostic and non-prescriptive.
+Its purpose is not to fix verification practices in advance, but to offer a stable coordination layer between legal text and assessment practice that can be extended, refined, and instantiated by different actors, including sectoral initiatives and EU-funded testing infrastructures.
+In this sense, the contribution of this work lies in establishing a shared operational grammar for AI Act compliance verification, one that enables consistency without rigidity and innovation without fragmentation.
+Acknowledgement
+This work is supported by the EU project Citcom.AI and Vinnova INTERSTICE project (reference number: 2024-00661). This work is also supported by the Swedish AI Factory and the Luxembourg AI Factory.

### v3

@@
 1. Introduction
-Progress in AI has created substantial opportunities while also introducing significant technical, organisational, and regulatory challenges, particularly as AI systems are increasingly deployed in critical domains where failures may have serious societal consequences (bengio2024managing; wang2022artificial).
-In response, regulatory efforts worldwide are intensifying to promote trustworthiness in AI development and deployment (huang2024; smuha2021race; novelli2024taking).
-Within this context, the EU Artificial Intelligence Act (AI Act) (eu_ai_act_2024) represents one of the most comprehensive regulatory responses to date, establishing a risk-based framework that mandates ex ante verification and ongoing oversight for high-risk AI systems.
+Progress in AI has created substantial opportunities while also introducing significant technical, organisational, and regulatory challenges, particularly as AI systems are increasingly deployed in critical domains where failures may have serious societal consequences (Bengio et al., 2024; Wang and Chung, 2022).
+In response, regulatory efforts worldwide are intensifying to promote trustworthiness in AI development and deployment (Huang et al., 2024; Smuha, 2021; Novelli et al., 2024).
+Within this context, the EU Artificial Intelligence Act (AI Act) (40) represents one of the most comprehensive regulatory responses to date, establishing a risk-based framework that mandates ex ante verification and ongoing oversight for high-risk AI systems.
 Despite this progress, a persistent gap remains between high-level legal requirements and the concrete verification activities needed to demonstrate compliance in practice.
@@
 These challenges are compounded by three interrelated forms of uncertainty.
-Interpretive uncertainty concerns the contextual meaning and scope of legal obligations; operational uncertainty relates to how such obligations can be instantiated across heterogeneous development and deployment practices; and procedural uncertainty arises from evolving standards, guidance documents, and the gradual designation of notified bodies (deloitte2024ai; dlapiper2025pause; draghi2024competitiveness; arnal2024implementation; lewis2025regulatorylearning; regulatinguncertainty2025).
-These uncertainties are further amplified by the emergent behaviour of modern generative AI systems, whose properties arise from complex interactions rather than explicit programming (holtzmanGenerativeModelsComplex2025; koltLessonsComplexSystems2025a).
-As a horizontal regulation relying on harmonised standards, post-market monitoring, and iterative guidance, the AI Act embodies a model of regulatory learning in which compliance expectations are progressively refined through implementation experience and verification outcomes (zakiConceptualisingOrganisationalPolicy2025; lewis2025regulatorylearning; schrepelAdaptiveRegulation2025).
+Interpretive uncertainty concerns the contextual meaning and scope of legal obligations; operational uncertainty relates to how such obligations can be instantiated across heterogeneous development and deployment practices; and procedural uncertainty arises from evolving standards, guidance documents, and the gradual designation of notified bodies (Deloitte, 2024; Piper, 2025; Draghi, 2024; Arnal, 2024; Lewis et al., 2025; unknown, 2025).
+These uncertainties are further amplified by the emergent behaviour of modern generative AI systems, whose properties arise from complex interactions rather than explicit programming (Holtzman et al., 2025; Kolt et al., 2025).
+As a horizontal regulation relying on harmonised standards, post-market monitoring, and iterative guidance, the AI Act embodies a model of regulatory learning in which compliance expectations are progressively refined through implementation experience and verification outcomes (Zaki, 2025; Lewis et al., 2025; Schrepel, 2025; Mowla and Kabir, 2024).
 Such learning depends on the availability of structured, comparable, and traceable verification evidence, which remains fragmented across actors, disciplines, and use cases.
@@
 Existing initiatives provide partial foundations for AI verification, but none offers a complete and operational mapping between legal requirements and verification activities.
-The NIST AI Risk Management Framework (NIST_AI_RMF_2025) provides high-level guidance but is not aligned with EU-specific legal obligations, while ISO/IEC standards such as ISO/IEC 42001 (iso_iec_42001_2023) establish process-oriented management requirements without specifying AI Act–specific verification procedures.
-European standardisation efforts within CEN–CENELEC JTC 21 (CEN-CENELEC_JTC21_2025) aim to develop harmonised standards granting presumption of conformity under Article 40, but this work remains ongoing.
+The NIST AI Risk Management Framework (National Institute of Standards and Technology (NIST), 2025) provides high-level guidance but is not aligned with EU-specific legal obligations, while ISO/IEC standards such as ISO/IEC 42001 (International Organization for Standardization and International Electrotechnical Commission, 2023) establish process-oriented management requirements without specifying AI Act–specific verification procedures.
+European standardisation efforts within CEN–CENELEC JTC 21 (CEN-CENELEC Joint Technical Committee 21, 2025) aim to develop harmonised standards granting presumption of conformity under Article 40, but this work remains ongoing.
 Against this background, the central contribution of this paper is a structured and operational mapping that translates high-level AI Act requirements into concrete verification activities applicable across the AI lifecycle.
@@
 While related efforts exist (e.g.
-(hernandez2025open)), the proposed mapping provides a higher level of granularity and operationalisability.
+(Hernandez et al., 2025)), the proposed mapping provides a higher level of granularity and operationalisability.
 It is not intended to be prescriptive or exhaustive, but to serve as a reusable reference that can evolve alongside regulatory guidance, standardisation efforts, and advances in evaluation techniques, thereby supporting more consistent conformity verification and evidence-based regulatory learning over time.
@@
 Given the heterogeneity and dispersion of these obligations across the regulation, an explicit structuring step is required to organise them into a coherent form suitable for systematic verification.
-For this purpose, we use the seven principles of Trustworthy AI defined by the European Commission’s High-Level Expert Group in the Ethics Guidelines for Trustworthy AI (eu_trustworthy_ai_2019) as an organising abstraction.
+For this purpose, we use the seven principles of Trustworthy AI defined by the European Commission’s High-Level Expert Group in the Ethics Guidelines for Trustworthy AI (High-Level Expert Group on Artificial Intelligence, 2019) as an organising abstraction.
 Although these principles predate the AI Act and are not legally binding, they have strongly influenced the development of European AI governance and are reflected throughout the regulation, both as explicit obligations (e.g.
@@
 These sources are predominantly non-binding, complemented where appropriate by binding instruments whose established operational mechanisms are widely reused in compliance and assurance activities.
-In this respect, the GDPR (GDPR2016) is included not as a source of normative extension of the AI Act, but because it provides mature and widely operationalised mechanisms—such as risk-based assessment, documentation practices, and accountability procedures—that directly inform how several AI Act obligations are interpreted and verified in practice.
+In this respect, the GDPR (European Union, 2016) is included not as a source of normative extension of the AI Act, but because it provides mature and widely operationalised mechanisms—such as risk-based assessment, documentation practices, and accountability procedures—that directly inform how several AI Act obligations are interpreted and verified in practice.
 The AI Act explicitly operates in complementarity with existing data protection law, and high-risk AI systems in the European context frequently involve personal data processing throughout development, deployment, and monitoring.
@@
 Beyond this exception, operational grounding relies on non-binding but authoritative sources that provide reusable mechanisms for implementing and assessing regulatory obligations in practice.
-These include the EU General Purpose AI Code of Practice (gpai-code-practice-3rd), ISACA’s Advanced Audit in AI Official Review Manual (isaca2025aaia), UNESCO’s Recommendation on the Ethics of AI (unesco2021ethics), the Web Content Accessibility Guidelines (WCAG) 2.2 (wcag22), and a broad range of ISO and IEC standards relevant to AI governance, risk management, quality management, cybersecurity, and data quality (e.g.
-(ISO_IEC_22989_2022; ISO23894; ISO_IEC_TR_24027_2021; ISO_IEC_24029_2_2023; ISO_IEC_24970_DIS_2025; ISO_IEC_27001_2022; ISO_IEC_27005_2022; ISO_IEC_27035_1_2023; ISO9001; ISO_31000_2018; ISO_IEC_31010_2019; iso_iec_42001_2023; ISO_IEC_5259_1_2024; ISO_8000_8_2015)).
+These include the EU General Purpose AI Code of Practice (Chairs and Vice‑Chairs of the General‑Purpose AI Code of Practice, 2025), ISACA’s Advanced Audit in AI Official Review Manual (ISACA, 2025), UNESCO’s Recommendation on the Ethics of AI (UNESCO, 2021), the Web Content Accessibility Guidelines (WCAG) 2.2 (World Wide Web Consortium (W3C), 2023), and a broad range of ISO and IEC standards relevant to AI governance, risk management, quality management, cybersecurity, and data quality (e.g.
+(27; 26; 28; 5; 3; 24; 25; 23; 31; 41; 42; International Organization for Standardization and International Electrotechnical Commission (2023); 4; 11)).
 These sources were selected because they translate high-level principles into implementable governance processes, controls, and audit practices, thereby bridging normative requirements and operational verification.
@@
 Highly cited and influential works in AI governance, robustness testing, risk management, and human–system interaction were considered (e.g.
-(Brundage2020; Basiri2016; Floridi2018; mundhenk2020; sheridan1992telerobotics; schuler1993participatory; CoE_AI_2020)).
+(Brundage and et al., 2020; Basiri et al., 2016; Floridi et al., 2018; Mundhenk et al., 2020; Sheridan, 1992; Schuler and Namioka, 1993; Leslie et al., 2020)).
 Selection criteria combined citation impact, venue relevance, and explicit connection to measurable or testable aspects of AI trustworthiness.
@@
 Numerous software and AI lifecycle models provide fine-grained decompositions of development and deployment phases (e.g.
-(de2022artificial; wang2021artificial)).
+(De Silva and Alahakoon, 2022; Wang et al., 2021)).
 However, such models vary significantly across standards, engineering practices, and organisational contexts, and their level of detail often presupposes technical expertise.
@@
 Requirement References Methods Type Target
-R1.1: Human in the loop Sheridan et al. (sheridan1992telerobotics) M1.1: Manual override — Confirm that a human operator can take control of the AI system at any time during operation. C P, FP
-Sheridan et al. (sheridan1992telerobotics) M1.2: Supervisory control — Assess whether human operators are able to supervise the AI system and intervene when necessary. C P, FP
+R1.1: Human in the loop Sheridan et al. (Sheridan, 1992) M1.1: Manual override — Confirm that a human operator can take control of the AI system at any time during operation. C P, FP
+Sheridan et al. (Sheridan, 1992) M1.2: Supervisory control — Assess whether human operators are able to supervise the AI system and intervene when necessary. C P, FP
 ISO/IEC 42001:2023 (Ann.A.3.2) M1.3: Strategic governance — Ensure that organisational oversight mechanisms for AI deployment are defined and documented. C P
 R1.2: User Autonomy GDPR (Art.7); Andreotta et al.
-(andreotta2022ai) M1.4: Informed consent — Check that users are provided with clear information about AI usage and are able to give or withhold consent.
+(Andreotta et al., 2022) M1.4: Informed consent — Check that users are provided with clear information about AI usage and are able to give or withhold consent.
 C FP
@@
 ISACA manual (§ 2.8.6) M2.2: Drift detection — Confirm that mechanisms are in place to monitor data distribution changes and evaluate behaviour under such shifts. T D, P
-Basiri et al. (Basiri2016) M2.3: Chaos engineering — Determine whether resilience has been tested through controlled fault injection in production-like environments. T P
+Basiri et al. (Basiri et al., 2016) M2.3: Chaos engineering — Determine whether resilience has been tested through controlled fault injection in production-like environments. T P
 ISACA manual (§ 2.9.4) M2.4: Fail-over architecture — Ensure redundant components exist to maintain functionality in case of failure. C P, FP
-R2.2: Robustness to Attacks Brundage et al. (Brundage2020); ISACA manual (§ 2.13.3) M2.5: Red teaming — Examine whether adversarial testing is conducted to uncover vulnerabilities. T P, FP
+R2.2: Robustness to Attacks Brundage et al.
+(Brundage and et al., 2020); ISACA manual (§ 2.13.3) M2.5: Red teaming — Examine whether adversarial testing is conducted to uncover vulnerabilities.
+T P, FP
 R2.3: Fallback & Fail-safe HLEG Guidelines M2.6: Human fallback — Check that the system includes a mechanism to transfer control to a human in the event of failure. C P, FP
@@
 T D, M, FP
-Mundhenk et al. (mundhenk2020) M4.2: Saliency maps — Check if visualisation tools highlight relevant input regions. T D, M, FP
+Mundhenk et al. (Mundhenk et al., 2020) M4.2: Saliency maps — Check if visualisation tools highlight relevant input regions. T D, M, FP
 R4.2: Model & Data Traceability GPAI CoP (Sec. Transparency) M4.3: Model cards — Ensure that standardised documentation is available and maintained. C M, P
@@
 R5.3: Inclusive Design WCAG 2.2 M5.4: Accessibility — Check that the system supports assistive technologies. C FP
-R5.4: Stakeholder Engagement Schuler et al. (schuler1993participatory) M5.5: Participatory design — Assess whether affected users and communities are involved in design and testing. C FP, P
+R5.4: Stakeholder Engagement Schuler et al. (Schuler and Namioka, 1993) M5.5: Participatory design — Assess whether affected users and communities are involved in design and testing. C FP, P
 3.6. R6: Societal and Environmental Well-being
@@
 R6.1: Environmental Impact UNESCO (Rec. 84); ISACA manual (§ 1.17.7) M6.1: Energy-use tracking — Verify that the system monitors computational energy consumption and carbon footprint. T P, FP
-R6.2: Ethical Alignment Leslie et al. (CoE_AI_2020) M6.2: Rights-based system — Assess whether system design aligns with fundamental rights. C P
+R6.2: Ethical Alignment Leslie et al. (Leslie et al., 2020) M6.2: Rights-based system — Assess whether system design aligns with fundamental rights. C P
 UNESCO (Rec. 58) M6.3: Independent ethics committee — Verify that an independent ethics committee or officer is appointed. C P
@@
 ISO/IEC 42001:2023 (Ann.B.8.3) M7.3: External reporting interfaces — Ensure that secure portals are available for third-party auditors. C P
-R7.3: Incident Reporting Floridi et al. (Floridi2018) M7.4: Redress systems — Check whether users can report issues or contest AI-driven outcomes. C FP, P
+R7.3: Incident Reporting Floridi et al. (Floridi et al., 2018) M7.4: Redress systems — Check whether users can report issues or contest AI-driven outcomes. C FP, P
 R7.4: Documentation Integrity ISO/IEC 42001:2023 (Ann.B.6.2.5); ISACA manual (§ 2.11.1) M7.5: Version-controlled records — Assess whether documentation is versioned and traceable over time. C P
@@
 While the mapping is generic and technology-agnostic, this example shows how it can structure assurance activities across the AI lifecycle, from design-time verification to validation and documentation.
-Table LABEL:tab:use_case presents the results of this mapping process; for the sake of brevity, it is reported only partially and serves as an illustrative example.
+Table LABEL:tab:use_case presents the results of this mapping process.
 The use case concerns an AI-based system for detecting cyberattacks within connected vehicles in an industrial setting at Scania.
@@
 The mapping enables structured and traceable verification of data handling, model behaviour, explainability, and documentation, providing evidence to support safety, cybersecurity, and regulatory requirements without intervening in system development.
+Table 12. Results of the mapping instantiation for the Scania use case.
 Req. Sub-req. Method Instantiation in the Use Case
@@
 This role is particularly relevant in a cross-border context, where divergent sandbox practices risk emerging in the absence of a common operational reference.
-European-funded initiatives such as AI Factories (eu_ai_factories) and Testing and Experimentation Facilities (TEFs) (eu_tefs) are expected to play an increasingly central role in supporting the testing and validation of AI systems, including in the context of regulatory sandboxes, as reflected in recent policy and implementation guidance (implementingAct).
+European-funded initiatives such as AI Factories (European Commission, 2025a) and Testing and Experimentation Facilities (TEFs) (45) are expected to play an increasingly central role in supporting the testing and validation of AI systems, including in the context of regulatory sandboxes, as reflected in recent policy and implementation guidance (European Commission, 2025b).
 In this setting, the availability of a shared mapping between legal requirements and verification activities becomes critical.